### v1

# MindGYM: Enhancing Vision-Language Models via Synthetic Self-Challenging Questions

Abstract
Large vision-language models (VLMs) face challenges in achieving robust, transferable reasoning abilities due to reliance on labor-intensive manual instruction datasets or computationally expensive self-supervised methods. To address these issues, we introduce MindGYM, a framework that enhances VLMs through synthetic self-challenging questions, consisting of three stages: (1) Seed Single-Hop Question Synthesis, generating cognitive questions across textual (e.g., logical deduction) and multimodal contexts (e.g., diagram-based queries) spanning eight semantic areas like ethical analysis; (2) Challenging Multi-Hop Question Synthesis, combining seed questions via diverse principles like bridging, visual-textual alignment, to create multi-step problems demanding deeper reasoning; and (3) Thinking-Induced Curriculum Fine-Tuning, a structured pipeline that progressively trains the model from scaffolded reasoning to standalone inference. By leveraging the model’s self-synthesis capability, MindGYM achieves high data efficiency (e.g., +16% gains on MathVision-Mini with only 400 samples), computational efficiency (reducing both training and inference costs), and robust generalization across tasks. Extensive evaluations on seven benchmarks demonstrate superior performance over strong baselines, with notable improvements (+15.77% win rates) in reasoning depth and breadth validated via GPT-based scoring. MindGYM underscores the viability of self-challenging for refining VLM capabilities while minimizing human intervention and resource demands. Code and data are released to advance multimodal reasoning research.
**footnotetext: Equal contributions.
1 Introduction
Large vision-language models (VLMs) have emerged as foundational tools for tasks requiring joint reasoning over visual and textual inputs [4, 30]. However, enabling robust performance with transferable and efficient reasoning abilities with high-quality data remains challenging [6, 8]. Manually curated instruction datasets like OK-VQA [25] and ScienceQA [22] are labor-intensive to scale, while self-supervised synthetics methods such as MMInstruct [21] and MMEvol [24] suffer from limited generalization cross modality and tasks, failing to tightly and logically couple visual and textual semantics. Meanwhile, reasoning enhancement methods—such as reinforcement learning (RL) [1] or iterative prompting [19]—incur prohibitive computational costs, limiting their practicality.
To address these challenges, we propose MindGYM, a new framework illustrated in Figure 1 that enhances VLMs via synthetic self-challenging questions with three core stages: (1). Seed Single-Hop Question Synthesis: Generates cognitive questions across text-only (e.g., logical deduction) and multimodal contexts (e.g., scientific diagrams), spanning eight semantic topics like causal reasoning and ethical analysis. (2). Challenging Multi-Hop Question Synthesis: Combines seed questions using cognitive operators—textual strategies (Bridging, Comparison, Temporal) and multimodal reasoning types (Visual-Textual Alignment, Spatial Reasoning, etc.)—to create layered, multi-step problems demanding deeper cognitive engagement. (3). Thinking-Induced Curriculum Fine-Tuning: A structured training pipeline progresses from scaffolded reasoning (e.g., answer prediction with explicit thought paths) to standalone inference, embedding advanced reasoning capacities into model parameters. All these processes are operated on the model to be tuned, via cognitive data self-synthesis and self-challenging training.
Extensive empirical evaluations across 7 benchmarks demonstrate that MindGYM consistently outperforms prior methods on both text-based and multimodal scenarios, achieving +16% gains on challenging tasks like MathVision-Mini with merely 400 synthesis samples, while reducing computational costs in both training and inference time. Notably, MindGYM gains robust generalization across different synthesis sources, and enhanced reasoning depth and breadth via GPT-based scoring analysis over state-of-the-art (SOTA) competitors (+15.77% win rates), enabling the model to handle open-ended tasks requiring multi-modal, high-order cognition.
Our main contributions can be summarized as follows:
  • We propose a new cognitive-grounded synthetic data synthesis pipeline, that systematically generates complex questions considering thinking depth, thinking width, and cross-modality.
  • We show the feasibility of self-challenging evolution for VLMs, with high data-efficiency and progressive training that absorbs explicit human-like reasoning patterns.
  • We verify MindGYM’s effectiveness, practical efficiency, and contextual adaptability with extensive experiments. Our code and data are released at https://github.com/modelscope/data-juicer/tree/MindGYM to foster future progress in vision-language reasoning.
Figure 1: The proposed MindGYM framework architecture consists of three core stages: ① Seed Single-Hop Question Synthesis, ② Challenging Multi-Hop Question Synthesis, and ③ Thinking-Induced Curriculum Fine-Tuning. The process begins with cognitive-aware self-synthesis across text (e.g., mathematical/biological questions) and multimodal (text+image) contexts. Seed questions are then combined via text-based strategies (Bridging, Comparison, Temporal) and multimodal reasoning types (Visual-Textual Alignment, Spatial Reasoning, Causal Inference, Contextual Synthesis) to form self-challenging questions. Finally, a curriculum-driven fine-tuning phase injects structured cognitive competence into the model through masked-task learning. Color-coded regions and arrows illustrate the hierarchical progression from simple reasoning to advanced problem-solving, emphasizing conceptual depth and modal integration.
2 Related Works
Instruction Data Construction for VLMs.
Large vision-language models (VLMs) critically depend on high-quality instruction-tuning datasets to achieve robust performance [imgdiff, 7]. Early efforts, such as OK-VQA [25] and ScienceQA [22], rely on human-annotated data to ensure accuracy and relevance. However, these manual approaches are resource-intensive and difficult to scale. To address this limitation, recent works have explored self-instruct paradigms that automate dataset creation. While text-centric methods like [39, 34, 42] emphasize diversity and complexity in textual instructions, their extension to multimodal domains remains challenging. Multimodal self-instruction frameworks, such as MMInstruct [21] and MMEvol [24], attempt to bridge this gap by iteratively generating image-text pairs. Nevertheless, these methods often suffer from modality inconsistency, where generated textual instructions lack tight visual grounding or rely on shallow cross-modal alignments (e.g., generic image descriptions). This issue arises due to an over-reliance on LLMs for text generation, which lack explicit mechanisms for visual reasoning. Consequently, the resulting datasets fail to challenge VLMs on fine-grained vision-language understanding tasks.
Reasoning-Aware Question Synthesis.
Synthetic question generation has shown success in domains like mathematical reasoning [20, 36], where structured problem spaces enable systematic scaling of complexity [32] and skill decomposition [3]. However, extending these techniques to multimodal domains poses unique challenges due to three key factors: (1) Heterogeneous Semantics: Vision-language reasoning requires handling unstructured inputs with diverse modalities, unlike the formal grammars available in mathematical problems. (2) Cross-Modal Synergy: Current methods focus on single-modality skill progression rather than fostering interactions between modalities. (3) Open-Ended Challenges: Template-driven approaches, such as Visual Program Induction [12], struggle to generate open-ended reasoning tasks that demand higher-order cognitive skills. As a result, existing methods exhibit limited generalizability to complex vision-language tasks, underscoring the need for more flexible and domain-agnostic synthesis frameworks.
Reasoning-Aware Self-Improvement.
Enhancing reasoning capabilities in VLMs often incurs significant computational overhead. RL-based approaches [1] require extensive policy optimization over large-scale reasoning trajectories [14], making them computationally expensive. Iterative refinement techniques [19] further exacerbate costs by introducing multi-stage processing pipelines. Even parameter-efficient methods, such as Chain-of-Thought prompting [40], suffer from latency penalties due to lengthy sequential decoding. Recent advances in test-time scaling [26] demonstrate promising performance gains but at the expense of increased inference latency. These limitations highlight the need for more efficient strategies to improve reasoning without compromising computational feasibility.
Our Position.
MindGYM addresses these three critical limitations of prior art: (1) Modality-inconsistency. With multiple cognitive relationship enhancement in multi-hop questions, we enforce tight semantic coupling in vision-language concepts. (2) Generalization. By adopting domain-agnostic prompts, we explicitly incorporate eight vision meta-topics beyond math, and adaptively increase visual grounding and linguistic complexity on the synthesis data. (3) Computational efficiency. By injecting complex, self-challenging questions directly into the training data, we encourage models through curriculum fine-tuning to perform deep and wide reasoning within a single inference step, eliminating the need for iterative adjustments.
3 The Proposed MindGYM
In this work, we aim to embed advanced reasoning capabilities into VLMs through synthesized fine-tuning datasets. For large models, conventional single-hop question-answer pairs are insufficient to stimulate reasoning in both depth and breadth. To address this, we propose a progressive self-synthesis framework, outlined in Figure 1, operating through three interconnected stages:
(1) Seed Question Synthesis: Starting with a base VLM $\pi$ (e.g., Qwen-2.5-VL [4]), the model self-generates logically related single-hop questions $\{q^{1}_{i},\ldots,q^{k}_{i}\}_{i\in N}$ to explore diverse reasoning methodologies and topics. Here, $k$ denotes the number of seed questions per context and $N$ represents the number of synthesized data.
(2) Self-Challenging Question Synthesis: Using $\{q^{1}_{i},\ldots,q^{k}_{i}\}$ as inputs, the model constructs structured complex questions $\{(Q_{i},A_{i},T_{i})\}_{i\in N}$ via multi-hop combinations. Each $Q_{i}$ is formed as $Q_{i}=q^{1}_{i}\oplus\ldots\oplus q^{k}_{i}$, where $\oplus$ represents cognitive operators (e.g., composition, contradiction, abstraction) that create hierarchical reasoning structures. The model concurrently generates explicit thinking trajectories $T_{i}$, which documents the rationale behind $Q_{i}$ and $A_{i}$.
(3) Self-Challenging Cognitive Training: The model is fine-tuned on four structured datasets $\{D^{j}_{\text{train}}\}_{j=1}^{4}$ through a curriculum pipeline. Training focuses on predicting masked components across questions $(Q_{i})$, answers $(A_{i})$, and their associated thought paths $(T_{i})$, thereby cultivating deeper and broader reasoning capacity. This adversarial questioning approach systematically improves the model’s ability to navigate ambiguous, multi-step reasoning tasks.
The proposed framework enhances generalization by exposing models to diverse cognitive challenges while injecting interpretable thinking patterns that mirror human-like problem-solving dynamics into the model parameters. Detailed design and implementations are provided in subsequent subsections.
3.1 Seed Single-Hop Question Synthesis
3.1.1 Diverse Thinking Methodologies
Cognitive-Aware Synthesis Templates.
We adopt a synthesis template that injects thinking elements, consisting of three parts: cognitive setting, background conditions, and synthesis goals: $P=\{P_{cog},P_{context},P_{task}\}$. $P_{context}$ and $P_{task}$ have different instantiations in different synthesis scenarios (single-hop, multi-hop QA, text, multimodal) in our method  [^1] The full prompt is presented in the Appendix B., while $P_{cog}$ remains unchanged to encourage the model to consistently produce responses that are strongly related to thinking.
We first describe the single-hop question scenario, with other tasks detailed later. Based on the given $P_{context}$, $P_{task}$ requires $\pi$ to generate $k$ logically connected seed single-hop questions $q^{1}_{i},...q^{k}_{i}$, at the same time, the model will be explicitly noted that these seed questions will be used later as more complex multi-hop combinations (the specific combination types will be detailed in later Sec.3.2).
Adaptive Stream of Consciousness.
We draw inspiration from a thinking protocol named Thinking-Claude [44] and adapt it to shape $P_{cog}$ as a collection of many diverse thinking processes, (e.g., making and challenging hypotheses, exploring and balancing different perspectives). This makes the model mimic the human thinking process in a layered, context-aware style, aiming for a broader coverage in thinking methodologies. Given different $P_{context}$ and $P_{task}$, $\pi$ dynamically adjusts thinking dimensions, such as problem depth, interpretation angles, and validation rigor, and explicitly expose models’ “minds” dealing with complex scenarios. This adaptive exploration not only encourages the model to incorporate diverse reasoning paths for generalizing to reasoning tasks, but also injects structured variability into training data, countering the memorization biases in conventional pipelines. Empirical validation of $P_{cog}$ is presented in Section 4.3.
3.1.2 Diverse Thinking Objects
Meta-Topics.
We consider both text-only and text-image QA synthesis. Here we first introduce the $P_{context}$ for the text-only case. Building on the cognitive architecture theory [28] and the multi-compartmental knowledge representation [33], we establish eight orthogonal meta-topics spanning four foundational reasoning dimensions: quantitative, causal, temporal, and social-ethical. Unlike prior works focused on limited domain combinations [2], our approach emphasizes broad cognitive coverage:
  • Mathematical Reasoning: Solving quantitative problems using numbers and formulas.
  • Scientific Knowledge: Understanding natural laws and scientific principles.
  • Logical Deduction: Forming conclusions through logical progression from premises.
  • Technical Procedures: Following step-by-step instructions for practical tasks.
  • Historical Events: Analyzing past events and their consequences.
  • Ethical Considerations: Evaluating decisions through moral frameworks.
  • Economic Principles: Studying resource allocation and valuation.
  • Psychological Insights: Exploring human behavior and cognition.
These eight topics are selected for two core reasons: (1) alignment with widely studied reasoning domains (e.g., mathematics, logic) [17], and (2) extension to synergize with VLMs’ capabilities and potential [30], balancing foundational reasoning with subsequent self-challenges.
By polling one of these meta-topics and their descriptions, we require $\pi$ to self-generate background document as $P_{context}$, then self-generate $k=5$ seed single-hop questions per inference. Empirical results in Section 4.3 demonstrate that this topic-balanced sampling strategy enhances thinking breadth and significantly improves end-to-end model performance compared to unbalanced alternatives. This process allows $\pi$ to explore maximal topic diversity, leveraging its internal knowledge to identify domains and skills requiring advanced reasoning.
Anchoring Image Sources.
For the $P_{context}$ component in text-image QA synthesis, we use existing images rather than generating images from scratch (as in the text-only case). This is because the existing SOTA VLMs are mostly only image-to-text generation (i.e., image understanding) and fail to do text-to-image generation at the same time. This decision arises because current SOTA VLMs predominantly specialize in image-to-text generation (i.e., image understanding) and lack dual capabilities for text-to-image generation constrained by their architectures. Without anchored visual references, the model risks hallucinatory reasoning about image elements (e.g., referencing “growth curves” might lead to incorrect inferences about unrelated scenarios, such as spatial-based reasoning tasks).
Specifically, we randomly sample image-text pairs from scientific knowledge and general-domain QA datasets – ScienceQA [22] and OKVQA [25] – to serve as background context ($P_{\text{context}}$) for $\pi$. We emphasize alignment between visual and textual concepts within the generation instruction $P_{\text{task}}$, prompting $\pi$ to self-generate $k=5$ seed questions per inference. This enhances harmonized interactions between image and text modalities during synthesis.
Reject Sampling for Diversity.
To enhance diversity in seed question generation, MindGYM further applies semantic vectorization to textual components of data in the synthesis pool. We use an empirically determined cosine similarity threshold to evaluate the overlap between newly generated and existing samples. If similarity exceeds this threshold, the system initiates a regeneration cycle to eliminate semantically redundant content, continuing until the seed question pool reaches the target size $N$. This mechanism widens the breadth of reasoning in synthetic data, better preparing it for diverse downstream scenarios.
Unlike prior methods concentrating on domain-specific reasoning (e.g., mathematical tasks), MindGYM expands reasoning scope across three dimensions: cognitive-aware thinking methodologies, image-text interactions, and semantic redundancy filtering. In Section 4.2, we empirically validate MindGYM ’s robustness and effectiveness across varied synthesis sources—including model’s internal knowledge, multilingual contexts, and anchored visual content.
3.2 Challenging Multi-Hop QA Synthesis
3.2.1 Cognitive Combination Principles
Synthesis Templates for In-depth Thinking.
Recall that in Section 3.1, we generate $k$ single-hop seed questions and proactively inform $\pi$ that these questions will be used for future cognitive combinations. Here, we explicitly specify some cognitive combination relationships into the synthesis requirement $P_{task}$ for multi-hop QAs, and let $\pi$ to self-challenging by figuring out how to combine the generated $\{q^{1}_{i},...q^{k}_{i}\}_{i\in N}$ into more complex questions $\{Q_{i}\}_{i\in N}$ conditioned on aforementioned $P_{cog}$ and $P_{context}$, such that the results is logically right, cognitively considered and match the most suitable relationships among the given ones.
Building on the single-hop seed questions generated in Section 3.1, we advance to multi-hop question synthesis by explicitly defining cognitive combination relationships within the task specification $P_{\text{task}}$. During this phase, the model $\pi$ is prompted to synthesize more complex questions $\{Q_{i}\}_{i\in N}$ by logically combining the pre-generated seed questions $\{q^{j}_{i}\}_{i=1,...,N;j=1,...,k}$ under constraints from $P_{\text{cog}}$ (cognitive framework) and $P_{\text{context}}$ (factual reference). This process aims to make $\{Q_{i}\}_{i\in N}$ logically valid, cognitively grounded, and optimally match the most relevant relationships from our predefined taxonomy.
Textual Cognitive Combination Types.
For $P_{\text{context}}$ containing only textual content, we define three cognitive combination forms after initial seed generation:
  • Comparison: Requires contrasting two or more concepts or scenarios.
  • Bridging: Connects disparate concepts through intermediate reasoning steps.
  • Temporal: Involves analyzing events across chronological contexts.
Multimodal Cognitive Combination Types.
For $P_{\text{context}}$ containing both text and images, we introduce four multimodal combination types:
  • Visual-Textual Alignment: Ensures consistency between visual data (e.g., diagrams, charts) and textual descriptions.
  • Spatial Reasoning: Addresses spatial relationships or geometric arrangements depicted in visual content.
  • Causal Inference: Requires deducing cause-effect relationships from combined textual and visual inputs.
  • Contextual Synthesis: Demands integration of information across modalities (e.g., text, images, graphs) to form unified conclusions.
3.2.2 Cognitive Manifestation
Structured Cognitive Processes & Reasoning Units.
Beyond specifying cognitive combination relationships in the task requirement $P_{\text{task}}$, we explicitly require $\pi$ to articulate its intermediate thought process $T_{i}$ leading to the final question $Q_{i}$. This support $T_{i}$ links the seed questions $\{q^{j}_{i}\}_{i\in N}$ into a logically coherent reasoning pathway, explicitly connecting steps to ensure interpretability and contextual consistency.
To balance cognitive context relevance and inference efficiency, $\pi$ is prompted to jointly generate $Q_{i}$, $A_{i}$, and $T_{i}$ in one inference step. However, raw responses may include fragmented or overly lengthy content, so we further prompt $\pi$ to segment outputs into distinct components: formal question-answer pairs $(Q_{i},A_{i})$ and their supporting rationale $T_{i}$. This results in modular cognitive units of varying granularity, enabling more flexible refinement of $\pi$’s reasoning abilities in subsequent training (more details are in later Section 3.3).
Figure 2: An end-to-end example of MindGYM. For multimodal data, we generate five seed questions first, and make the model self-challenging itself via synthesizing multi-hop questions and multi-hop answers while preserving its internal thinking process.
Cognitive Information Processing Dynamics.
Note that the synthesis framework explicitly involves $\pi$ processing both novel and existing knowledge during inference. By reintroducing context elements from $P_{\text{context}}$, accessing seed questions $\{q^{j}_{i}\}_{i\in N}$, and addressing advanced combination requirements in $P_{\text{task}}$, the model is incentivized to: (1). Explore diverse reasoning patterns embedded in $P_{\text{cog}}$, (2). Re-examine context details with renewed emphasis, and (3). Construct multi-hop answers using enriched relational constraints. This process yields more focused, nuanced questions $Q_{i}$ while eliminating reasoning shortcuts, forcing $\pi$ to engage in deeper inferential analysis.
Figure 2 illustrates an example in which the original $P_{\text{context}}$ (including image and textual inputs) is first used to synthesize seed questions (e.g., “What ingredients did Kenji use?” or “What vitamins are in broccoli?”), as detailed in Section 3.1). These are then combined via the multi-hop synthesis protocol (outlined in Section 3.2) into more challenging questions that require cross-contextual reasoning (e.g., “How do Kenji’s lunch ingredients compare with Tracy’s in terms of vitamin content?” paired with bridging and temporal relationships).
Critically, whereas the original QA focused on limited details (e.g., Kenji’s tomatoes and Tracy’s broccoli), our synthesized questions expand reasoning depth by incorporating broader contextual elements: sandwich compositions, citrus fruits, root vegetables, etc. The resulting multi-hop QAs demonstrate richer interconnections, broader scope, and higher cognitive demands on the solver.
3.3 Thinking-Induced Curriculum Fine-Tuning
Training Objectives.
Given the self-synthesized dataset containing triples $\{(Q_{i},A_{i},T_{i})\}_{i\in N}$, we propose a “MindGym”-style training flow to progressively integrate cognitive guidance into the model $\pi$ through mixed masking strategies. This approach incrementally challenges $\pi$ by varying the exposure to question-answer pairs ($Q_{i}$, $A_{i}$), their supporting rationales ($T_{i}$), and masked input-output configurations.
For recap, $Q_{i}$ represents the complex multi-hop question, $A_{i}$ denotes the reference answer with structured reasoning, and $T_{i}$ records the detailed step-by-step cognitive trajectory. We construct four training datasets with distinct masking objectives:
$\mathcal{D}^{1}_{\text{train}}=\{(Q_{i},T_{i}\rightarrow A_{i})\}$ employs the masked input structure $\langle Q_{i},T_{i},[\mathbf{\text{MASK}}]\rangle$. The model predicts $A_{i}$ using $Q_{i}$ and explicit thinking guidance $T_{i}$, and the loss is defined as:
$\mathcal{L}_{1}=-\mathbb{E}_{(Q,T,A)\sim\mathcal{D}^{1}_{\text{train}}}\log\pi_{\theta}(A|Q,T).$ (1)
$\mathcal{D}^{2}_{\text{train}}=\{(Q_{i},A_{i}\rightarrow T_{i})\}$ uses the structure $\langle Q_{i},[\mathbf{\text{MASK}}],A_{i}\rangle$. The objective is to recover the hidden thinking path $T_{i}$, with loss:
$\mathcal{L}_{2}=-\mathbb{E}_{(Q,A,T)\sim\mathcal{D}^{2}_{\text{train}}}\log\pi_{\theta}(T|Q,A).$ (2)
$\mathcal{D}^{3}_{\text{train}}=\{Q_{i}\rightarrow(A_{i},T_{i})\}$ follows $\langle Q_{i},[\mathbf{\text{MASK}}]\rangle$, requiring simultaneous generation of $A_{i}$ and $T_{i}$:
$\mathcal{L}_{3}=-\mathbb{E}_{(Q,A,T)\sim\mathcal{D}^{3}_{\text{train}}}\log\pi_{\theta}(A,T|Q).$ (3)
$\mathcal{D}^{4}_{\text{train}}=\{Q_{i}\rightarrow A_{i}\}$ uses the same masking strategy as $\mathcal{D}_{train}^{3}$, with input $\langle Q_{i},[\mathbf{\text{MASK}}]\rangle$. This mimics standard supervised fine-tuning practice:
$\mathcal{L}_{4}=-\mathbb{E}_{(Q,A)\sim\mathcal{D}^{4}_{\text{train}}}\log\pi_{\theta}(A|Q).$ (4)
The curriculum progresses through these datasets sequentially: $\mathcal{D}^{1}\rightarrow\mathcal{D}^{2}\rightarrow\mathcal{D}^{3}\rightarrow\mathcal{D}^{4}$. This mirrors a guided learning pathway where $\pi$ first reproduces answers with external cognitive scaffolding ($\mathcal{D}^{1}$), then deduces thinking paths without full guidance ($\mathcal{D}^{2}$), gradually internalizes end-to-end reasoning ($\mathcal{D}^{3}$), and finally masters standalone Q-to-A mappings ($\mathcal{D}^{4}$). Empirical validation in Sec. 4.3 shows this strategy outperforms non-curriculum mixed training.
Discussion on Efficiency and Cost.
The MindGYM framework enhances $\pi$’s reasoning capabilities through $N\times 4$ synthesis inferences across stages ($P_{\text{context}}$, seed questions, $Q_{i}$, and $\mathcal{D}_{\text{train}}$). Despite this incremental setup, inference efficiency is maintained through amortized reasoning costs during training. A single inference suffices to achieve high model performance, as shown empirically in Section 4.3. Unlike test-time scaling approaches [38, 18], we move cognitive enhancement costs to the training phase. Compared to data-intensive methods (e.g., reinforcement learning) requiring extensive trial-and-error [19], our method maximizes cognitive depth leveraging structured synthetic data without resource-heavy exploration phases.
Datasets Multimodal Eval Text-Only Eval Avg
MMStar MathVista MathVision MM-Avg GSM8K MATH GPQA Text-Avg
Baslines Raw 64.00 69.30 24.67 52.66 83.62 67.60 31.83 61.02 56.84
LIMO (817) [43] 63.93 70.20 25.9 53.37 84.08 67.80 30.58 60.82 57.10
Open-O1 (77k) [35] 58.93 61.40 17.43 45.92 77.94 59.00 28.19 55.04 50.48
Open-O1 (400) [35] 63.87 69.70 23.66 53.08 84.31 69.00 29.7 61.00 57.04
MMEvol-SciQA (106) [24] 63.93 69.50 23.36 52.26 83.40 65.60 31.83 60.28 56.27
MMEvol-DVQA (4k) [24] 62.93 67.40 24.01 51.45 83.85 65.80 33.08 60.91 56.18
Text MindGYM (400) 64.33 70.30 28.62 54.42 84.08 68.40 31.33 61.27 57.84
Image MindGYM-OKVQA (400) 63.60 69.20 28.29 53.70 84.00 68.20 31.08 61.09 57.40
MindGYM-SciQA (400) 64.20 69.80 25.33 53.11 83.93 69.00 31.45 61.46 57.29
Table 1: Performance of Qwen2.5-VL-7B on various evaluation benchmarks. The number following the dataset name indicates its size in terms of the number of samples used for training.
4 Experiments
To comprehensively evaluate the performance of our proposed synthetic data method, we conduct extensive experimental validation on multiple representative text and multimodal evaluation sets, comparing it with several competitive data synthesis methods and datasets. Full descriptions of datasets, benchmarks, implementation details, and baselines are provided in Appendix A.
4.1 Experimental Settings
Datasets.
We conduct multi-modality data synthesis across textual and image contexts. For textual synthesis, data generation occurs exclusively through self-contained knowledge extraction from the base model, eliminating external data dependencies. Multimodal synthesis utilizes established scientific and reasoning corpora as visual primitives, with visual grounding from ScienceQA [22] and OK-VQA [25].
Benchmarks.
To systematically assess MindGYM on VLMs across reasoning modalities, we incorporate:
  • Multimodal evaluation sets include MMStar [9], MathVista-Mini [23], and MathVision-Mini [37], designed to comprehensively assess the reasoning abilities in the multimodal scenario.
  • Text-based evaluation sets comprise GSM8K [10], MATH [15], and GPQA [31], focused on evaluating reasoning and mathematical capabilities.
The MM-Avg and Text-Avg metrics serve as unified criteria for assessing multimodal and text-based reasoning capabilities respectively, whereas Avg quantifies comprehensive performance spanning both unimodal and multimodal reasoning domains.
Models & Implementation.
The data synthesis of MindGYM leverages VLMs’ inherent long-context understanding and reasoning capabilities. We implement Qwen2.5-VL-7B-Instruct [4] as our base VLM since alternative models (e.g., InternVL2.5-7B) exhibit inferior generation quality under identical configurations (detailed in Appendix D). To maintain linguistic consistency with Qwen2.5-VL-7B-Instruct and consider the information density of reasoning data, we utilize Chinese for data synthesis. All experiments follow consistent protocols: during text-only training, only LLM-layers update parameters, while image-text joint training additionally activates Aligner-layers, keeping ViT-layers permanently frozen.
Baselines.
As MindGYM generates self-contained adversarial training data without external sources, direct comparison with conventional data synthesis methods proves infeasible. We therefore select SOTA reasoning-oriented dataset works from two categories: (1) Multimodal reasoning using MMEvol’s [24] core subsets ScienceQA and DVQA, and (2) Text-based reasoning with LIMO [43], 817 curated logical chains, and Open-O1 [35], an SFT dataset for CoT activation as fundamental text reasoning baselines.
4.2 Main Results
To validate the efficacy of our methodology, we systematically evaluate the model across multiple benchmarks as documented in Table 1, with key insights summarized:
Universal Superiority.
MindGYM demonstrates substantial performance advantages through comprehensive average metric evaluation in MindGYM-Text with a 54.42 MM-Avg-score and 57.84 Avg-score, and notably achieving 28.62 on MathVision with a 16% improvement over the Raw model. The proposed method significantly outperforms all baselines where even the strongest competitor LIMO trails by 1.05 points in MM-Avg. An interesting phenomenon emerges among baselines that Open-O1 (77K) suffers severe degradation with 6.74-point MM-Avg decline and 29.3% MathVision collapse despite employing a large number of training samples, exposing the vulnerability of conventional scaling approaches.
Cross-modal Transfer.
MindGYM enables effective cross-modal reasoning capability sharing through bidirectional knowledge transfer. Remarkably, the text-specialized MindGYM-Text establishes new SOTA performance on MM-Avg, with the score of 54.42 that surpasses all baselines. Conversely, the multimodal-enhanced MindGYM-SciQA demonstrates unexpected supremacy in textual reasoning tasks, achieving an optimal Text-Avg score of 61.46. This bidirectional enhancement confirms the transferability of reasoning capabilities between textual and multimodal representations.
Generalizability & Robustness.
Consistent improvements emerge across diverse data sources. Our approach attains sub-optimal average performance when synthesizing data from both SciQA, showing an improvement of 0.45 over raw data, and OKVQA, with an improvement of 0.56 over raw data, demonstrating adaptability to various domain-specific reasoning tasks. This indicates that our method effectively enhances generalizability and robustness across diverse reasoning challenges.
Data Efficiency.
MindGYM actives model potential using only 400 synthesized samples, achieving a 28.62 MathVision score that surpasses baselines trained with 10$\times$-100$\times$ more data. This contrasts sharply with Open-O1’s performance collapse despite massive data scaling, indicating that MindGYM significantly enhances the model’s reasoning capabilities while also ensuring that the model’s performance is not detrimentally affected.
These findings effectively validate that our approach addresses the issues of modality inconsistency, generalization, and computational efficiency mentioned in Section 2.
4.3 Ablation Studies
To validate the efficacy of our approach, we conduct systematic module-wise utility analysis through targeted ablation studies.
GPT-based Thinking Quality Scoring.
Datasets Depth Breadth Avg
MindGYM/ Raw 10.2% 19.4% 14.8% $\uparrow$
MindGYM/ LIMO 1.63% 23.8% 12.7% $\uparrow$
MindGYM/ Open-O1 (400) 1.88% 23.7% 12.8% $\uparrow$
MindGYM/ MMEvol-DVQA 8.41% 37.0% 22.7% $\uparrow$
Overall 5.53% $\uparrow$ 26.0% $\uparrow$ 15.8% $\uparrow$
Table 2: The results of win rates in terms of relative improvements over baselines, using GPT-4 as a scorer.
We establish a two-dimensional evaluation framework leveraging GPT-4 as an expert scorer. The protocol assesses: (1) Reasoning Depth via derivation complexity, and (2) Reasoning Breadth by solution diversity. Each metric operates under a standardized scoring system, with detailed specifications provided in Appendix B.5. Comparative evaluations against four baselines are conducted on the MathVision dataset as a benchmark. The experimental outcomes documented in Table 2 demonstrate MindGYM’ consistent superiority across both evaluation axes, particularly achieving 15.8% average improvements in reasoning depth over the raw model and other competitive baselines.
Datasets MM-Avg Text-Avg Avg
Raw 52.66 61.02 56.84
MindGYM-Text 54.42 61.27 57.84
MindGYM-Text w/o SC 52.72 60.00 56.36 $\downarrow$
MindGYM-Text Syn-EN 52.77 60.30 56.53 $\downarrow$
MindGYM-Text w/o CF 52.20 60.83 56.51 $\downarrow$
MindGYM-Text with RB 52.20 60.83 56.51 $\downarrow$
MindGYM-OKVQA 53.70 61.09 57.40
MindGYM-OKVQA w/o SC 52.85 61.30 57.08 $\downarrow$
MindGYM-OKVQA Syn-EN 53.14 60.59 56.87 $\downarrow$
MindGYM-OKVQA with RB 53.33 60.60 56.97 $\downarrow$
MindGYM-SciQA 53.11 61.46 57.29
MindGYM-SciQA w/o SC 53.20 60.63 56.91 $\downarrow$
MindGYM-SciQA Syn-EN 52.25 60.88 56.56 $\downarrow$
MindGYM-SciQA with RB 53.46 60.27 56.87 $\downarrow$
Table 3: Ablation study results of removing stream of consciousness (w/o SC), utilizing English in data synthesis (Syn-EN), changing the order of Curriculum Fine-tuning (w/o CF) steps and Relation Balanced (with RB), complete results detailed in Appendix C.
Ablation of Stream of Consciousness.
We conduct the ablation study by removing the stream of consciousness module from our prompt while maintaining the same curriculum fine-tuning method for training and evaluation. The experimental results, as shown in Table 3, indicate that the removal of the stream of consciousness module leads to a decrease across all Avg scores, with reductions ranging from 0.32 to 1.48 points. This decline highlights the significance of the stream of consciousness module in our approach, underscoring its essential role in enhancing the depth of reasoning and overall performance of the model.
Curriculum Fine-tuning Order.
To verify the effectiveness of our Curriculum Fine-tuning module, we mix and shuffle the first three datasets ($\mathcal{D}^{1}_{\text{train}}$, $\mathcal{D}^{2}_{\text{train}}$, $\mathcal{D}^{3}_{\text{train}}$) and train the model on text data, subsequently using $\mathcal{D}^{4}_{\text{train}}$ to constrain the answers’ generation pattern. This experimental setup allows us to systematically examine whether disrupting the curriculum order impacts model performance. The experimental results shown in Table 3 under the w/o CF condition reveal a performance degradation of 1.33 points, highlighting the critical impact of the Curriculum Fine-tuning order on model performance.
Linguistic Difference.
Based on the selected Qwen series [4] model, we utilize Chinese as the language for data synthesis. Additionally, we explore the effects of using English for synthesis, as shown in Table 3 under the Syn-EN condition. The results indicate varying degrees of performance degradation when using English, with a decrease of up to 1.31 points in the MindGYM-Text setting. Beyond the inherent characteristics of Qwen, we hypothesize that the higher information density of Chinese makes it more beneficial for the model’s reasoning learning.
Relationship Distribution Balanced.
To verify the effectiveness of allowing the model to autonomously select relationship categories, we balance the number of entries generated for each category, totaling 400 entries. The experimental results shown in Table 3 under the with RB condition indicate that enforcing category balance leads to a decline up to 1.33 in overall model performance. This finding suggests that the autonomous selection of relationships by the model reflects its cognitive understanding.
5 Conclusion & Future Works
This work introduces MindGYM, a framework that enhances vision-language reasoning via synthetic self-challenging questions. By self-challenging multi-step reasoning data and employing structured curriculum training, MindGYM achieves strong performance gains with small data size and reduced computational costs, outperforming prior methods in reasoning depth and breadth.
Future directions include extending cognitive relationships to dynamic visual scenarios [46]; exploring the application of MindGYM in specific domains like medicine [27] and finance; developing adaptive scoring operators [6] for question complexity; and integrating agents [13] for data correctness verification in generalized RL environments.
Appendix A Implementation of Experiments
A.1 Utilized Seed Visual Datasets
For the multimodal domain, we leverage widely-adopted scientific and reasoning datasets as visual source material for question synthesis, specifically including ScienceQA and OK-VQA.
  • ScienceQA [22], the first large-scale multimodal dataset that annotates lectures and explanations for the answers.
  • OK-VQA [25], a dataset for visual question answering that requires methods which can draw upon outside knowledge to answer questions.
A.2 Benchmarks
To evaluate the models’ reasoning capabilities and performance in both multimodal and text comprehension, we adopt the methodology used by major advanced open-source VLM Qwen2.5-VL-7B [4], selecting the following widely utilized reasoning evaluation datasets. All evaluations were conducted on the ms-swift [45] platform.
  • Multimodal Benchmarks:
    • MMStar [9], an elite vision-indispensable multi-modal benchmark comprising 1,500 challenge samples meticulously selected by humans.
    • MathVista (Mini) [23], a benchmark designed to combine challenges from diverse mathematical and visual tasks, consists of 6,141 examples, derived from 28 existing multimodal datasets.
    • MathVision (Mini) [37], a meticulously curated collection of 3,040 high-quality mathematical problems with visual contexts sourced from real math competitions.
  • Text-based Benchmarks:
    • GSM8K [10], a dataset of high quality linguistically diverse grade school math word problems created by human problem writers, which takes between 2 and 8 steps to solve.
    • MATH [15], a dataset of challenging competition mathematics problems, each problem in MATH has a full step-by-step solution.
    • GPQA [31], stands for Graduate-Level Google-Proof Q&A Benchmark, a challenging dataset designed to evaluate the capabilities of LLMs and scalable oversight mechanisms.
Given the extensive evaluation tasks, employing the complete evaluation sets would incur substantial time expenditure. To expedite the evaluation process while ensuring fairness and accuracy, we adopt the $eval\_limit$ parameter from ms-swift, configuring GPQA to 300 samples, while retaining the original sample sizes for the remaining benchmark datasets. All experiments were conducted using this consistent setup to ensure the fairness of the experiments.
A.3 Training and Evaluation Details
Platform
We implement our approaches using PyTorch [29] v2.5.1, coupled with PEFT v0.14.0 and the Transformers library [41] v4.49.0. Experiments are conducted on a computing platform equipped with four NVIDIA A100 GPUs (40GB), with LLMs loaded as 16-bit floating-point numbers. The specific data-model development processes are completed in Data-Juicer Sandbox [8], via integration with the ms-swift [45] repository for training and evaluation, and the VLMEvalKit [11] repository for evaluation.
Baselines
Our method requires no external background knowledge for text generation and enhances reasoning capabilities through the model’s inherent visual understanding. While this approach differs fundamentally from typical data synthesis methodologies, we select representative reasoning-enhancing datasets for comparative analysis to validate the effectiveness of our approach in SFT for VLMs.
  • Multimodal Reasoning:
    • MMEvol[24]:A method addressing data quality limitations by generating complex and diverse image-text instruction datasets to improve VLMs. We consider two subsets of MMEvol: (1) For direct comparison, we utilize its ScienceQA synthetic data subset. (2) For comprehensive reasoning comparison, we adopt DVQA which primarily focusing on mathematical and logical reasoning tasks.
  • Text-Based Reasoning:
    • LIMO [43]: Challenges conventional assumptions in mathematical reasoning by demonstrating that models achieve superior performance with smaller quantities of high-quality training data. We reformat its ”question+solution+answer” structure into standard instruction pairs for training consistency.
    • Open-O1 [35]: An open-source initiative to replicate the reasoning capabilities of proprietary models through curated SFT data for CoT activation. The complete 77k-sample dataset trains both LLaMA and Qwen architectures. To evaluate data efficiency, we additionally create a 400-sample randomly subsampled version for performance validation on same scale.
Training Details
In our experimental setup, we employ Low-Rank Adaptation (LoRA) [16] adapters for the fine-tuning process, utilizing a LoRA-rank of 8 and a LoRA-alpha of 32. The learning rate was consistently maintained at $1\times 10^{-5}$ on Qwen2.5-VL-7B-Instruct. We utilize a batch size of 4 and set the maximum sequence length to 4096 tokens to accommodate the model’s reasoning capacity. To optimize the training process, a warmup ratio of 0.05 was applied and a validation ratio of 0.03 was used. The training was conducted over a single epoch, balancing computational efficiency with the need for effective model adaptation.
In the VLM training configuration, for pure textual information, we freeze both the Vision Transformer (ViT) and the Aligner while keeping the LLM layers active, to ensure that the image processing capabilities of the VLM remain unaffected. For image-text data, we adopt the post-tuning strategy of the Qwen2.5-VL series [4], where the ViT remains frozen while both the Aligner and LLM are activated to maximize multimodal information acquisition.
Evaluation Details
Following the evaluation protocols established in ms-swift [45] and VLMEvalKit [11], we maintained default configurations with pt as the inference backend. For multimodal tasks, we implemented VLMEvalKit as the evaluation backend, while employing the Native framework for text-only evaluation scenarios.
Appendix B Prompts
B.1 Stream of Consciousness
The protocol introduce in Section 3.1 is based on Thinking Claude, which is an innovative prompt engineering framework designed to enhance the reasoning capabilities of Claude AI models through simulated human-like cognitive processes. This paradigm-shifting approach employs a multi-stage cognitive architecture featuring chain-of-thought prompting, self-reflective mechanisms, and structured problem decomposition. We incorporate it into MindGYM’s data synthesis process to enhance the inference performance of the model [^2] https://github.com/richards199999/Thinking-Claude.
B.2 Prompts for Background Generation
On the text side, due to the low quality of sub-question generation without foundational documents, we create background documents for sub-question generation through iterative polling based on eight cognitive reasoning themes. On the image side, since text-to-image generation models are still underdeveloped and produce low-quality results, we directly utilize existing reasoning and trial-based QA datasets.
[…]
The document should be rich enough to support multi-hop reasoning in Chinese.
B.3 Prompts for Seed Single-Hop Question Synthesis
These two prompts are used to synthesize sub-problems, Prompt 2 is used to synthesize text sub-problems, and Prompt 3 is used to synthesize images sub-problems.
[…]
Based on the background document provided, generate up to 5 logically connected sub-questions. The relationship between these sub-questions should belong to a single category:
- Bridging: requiring connecting facts or pieces of information from the document.
- Comparison: involving comparing two or more elements described in the document.
- Temporal: requiring reasoning about the order or timing of events.
Clearly state the relationship category that links all sub-questions in Chinese:
[…]
Based on the provided image, original question, and original answer, generate up to 5 logically connected sub-questions. The relationship between these sub-questions should belong to one of the following categories:
- Visual-Textual Alignment: Requiring alignment between visual (e.g., images, charts) and textual information.
- Spatial Reasoning: Involving spatial relationships or geometric layouts.
- Causal Inference: Requiring reasoning about cause-and-effect relationships.
- Contextual Synthesis: Requiring synthesis of information across multiple modalities (e.g., text, images, charts).
Clearly state the relationship category that links all sub-questions in Chinese:
B.4 Prompts for Challenging Multi-Hop QA Synthesis
These two prompts are used to synthesize multi-hop problems, Prompt 4 is used to synthesize multi-hop problems of text, and Prompt 5 is used to synthesize multi-hop problems of images.
[…]
Synthesize a multi-hop question and its answer based on the above sub-questions and background document in Chinese. Please start Qwen thinking and return the thinking process:
[…]
The question should require reasoning across the sub-questions and synthesizing information from the image and original question.
Then, provide a detailed answer to the multi-hop question, ensuring it is consistent with the image and sub-questions.
Synthesize a multi-hop question and its answer based on the above sub-questions, original image, and original question in Chinese. Please start Qwen thinking and return the thinking process:
B.5 Prompts for GPT-based Scorer
This is a prompt scored with GPT4 to evaluate the scores of the two models in terms of depth and breadth.
[…]
As a mathematical problem solver, please strictly compare and analyze the predictions of the two models:
[Problem description]
{question}
[Original model prediction]
{pred_raw}
[Fine-tuned model prediction]
{pred_finetuned}
Evaluation requirements:
Give the scores of raw and finetuned respectively
1. Reasoning depth score (0-3 points):
- 3 points: multi-step derivation with verification
- 2 points: complete derivation steps
- 1 point: simple calculation steps
- 0 points: no derivation process
2. Reasoning breadth score (0-3):
- 3 points: Explores multiple valid methods/angles, justifies optimal choice
- 2 points: Mentions alternative approaches briefly but focuses on one
- 1 points: Suggests another method without analysis
- 0 points: Single approach with no alternatives considered
3. Comparative analysis: use bullet points to list the main improvements/degradations
Please return a JSON that strictly follows this format: ”accuracy_raw”: 0-2, ”accuracy_finetuned”: 0-2, ”depth_raw”: 0-3, ”depth_finetuned”: 0-3, ”comparison”: [”Point 1”, ”Point 2”]
Appendix C Complete Results including Ablation Study
We present the full experimental results in Table 4, comprehensively validating the effectiveness of our method compared to other baselines, as well as the robustness of our approach itself (ablation studies introduced in Section 4.3).
Datasets Multimodal Eval Text-Only Eval Avg
MMStar MathVista MathVision MM-Avg GSM8K MATH GPQA Text-Avg
Baslines Raw 64.00 69.30 24.67 52.66 83.62 67.60 31.83 61.02 56.84
LIMO (817) [43] 63.93 70.20 25.90 53.37 84.08 67.80 30.58 60.82 57.10
Open-O1 (77k) [35] 58.93 61.40 17.43 45.92 77.94 59.00 28.19 55.04 50.48
Open-O1 (400) [35] 63.87 69.70 23.66 53.08 84.31 69.00 29.70 61.00 57.04
MMEvol-SciQA (106) [24] 63.93 69.50 23.36 52.26 83.40 65.60 31.83 60.28 56.27
MMEvol-DVQA (4k) [24] 62.93 67.40 24.01 51.45 83.85 65.80 33.08 60.91 56.18
Text MindGYM 64.33 70.30 28.62 54.42 84.08 68.40 31.33 61.27 57.84
MindGYM w/o SC 63.60 69.90 24.67 52.72 83.83 66.60 29.57 60.00 56.36
MindGYM Syn-EN 63.93 69.70 24.67 52.77 84.00 67.20 29.70 60.30 56.53
MindGYM w/o CF 64.00 69.90 22.69 52.20 84.15 67.00 31.33 60.83 56.51
MindGYM w RB 64.00 70.10 22.70 52.20 84.15 67.00 31.33 60.83 56.51
Image MindGYM-OKVQA 63.60 69.20 28.29 53.70 84.00 68.20 31.08 61.09 57.40
MindGYM-OKVQA w/o SC 63.93 69.30 25.33 52.85 84.15 68.80 30.95 61.30 57.08
MindGYM-OKVQA Syn-EN 63.53 69.90 25.99 53.14 84.23 67.60 29.95 60.59 56.87
MindGYM-OKVQA w RB 63.80 71.20 25.00 53.33 84.00 66.60 31.20 60.60 56.97
MindGYM-SciQA 64.20 69.80 25.33 53.11 83.93 69.00 31.45 61.46 57.29
MindGYM-SciQA w/o SC 64.07 69.20 26.32 53.20 83.93 67.00 30.95 60.63 56.91
MindGYM-SciQA Syn-EN 63.73 69.00 24.01 52.25 83.55 68.00 31.08 60.88 56.56
MindGYM-SciQA w RB 64.20 70.20 25.99 53.46 83.85 65.00 31.96 60.27 56.87
Table 4: Complete ablation study results of removing stream of consciousness (w/o SC), utilizing English in data synthesis (Syn-EN), changing the order of Curriculum Fine-tuning (w/o CF) steps and Relation Balanced (with RB).
Appendix D Alternative VLMs Data Synthesis through MindGYM
As demonstrated in Listing 1, our analysis using InternVL2.5-8B [5] reveals its generated outputs tend to exhibit chaotic patterns. We hypothesize this stems from dual factors: (i) constrained input token limits restricting contextual comprehension, and (ii) inherent deficiencies in cognitive reasoning capabilities. These limitations necessitated its exclusion from our comparative VLM experiments.
Listing 1: Data Sample of MindGYM on InternVL2.5-8B.
1{
2 "background_document": "Background document for multi-hop reasoning.",
3 "reasoning_category": "Mathematical Reasoning",
4 "sub_questions": [
5 "What is the optimal pricing strategy considering cost inflation?",
6 "How will fixed and variable costs vary with seasonal demand?",
7 "What are the impacts of geopolitical changes on supply chain?",
8 "What role do regulatory changes play in pricing?",
9 "How will budget allocation differ among different product lines?"
10 ],
11 "relationship_category": "Comparison",
12 "multihop_question": "What steps need to be taken to optimize pricing strategies taking into account rising costs, shifting regulations, seasonal fluctuations, and differing product revenues?",
13 "multihop_answer": "To maximize profit under increased costs, companies should reassess fixed and variable expenses, implement strategic cost savings wherever feasible, monitor regulatory developments closely, review existing budgets, allocate resources accordingly, update pricing schedules, communicate revisions internally consistently, measure effectivity quarterly.",
[…]
15}

### v2

@@
-MindGYM: Enhancing Vision-Language Models via Synthetic Self-Challenging Questions
+MindGYM: What Matters in Question Synthesis for Thinking-Centric Fine-Tuning?
 Abstract
-Large vision-language models (VLMs) face challenges in achieving robust, transferable reasoning abilities due to reliance on labor-intensive manual instruction datasets or computationally expensive self-supervised methods.
-To address these issues, we introduce MindGYM, a framework that enhances VLMs through synthetic self-challenging questions, consisting of three stages: (1) Seed Single-Hop Question Synthesis, generating cognitive questions across textual (e.g., logical deduction) and multimodal contexts (e.g., diagram-based queries) spanning eight semantic areas like ethical analysis; (2) Challenging Multi-Hop Question Synthesis, combining seed questions via diverse principles like bridging, visual-textual alignment, to create multi-step problems demanding deeper reasoning; and (3) Thinking-Induced Curriculum Fine-Tuning, a structured pipeline that progressively trains the model from scaffolded reasoning to standalone inference.
-By leveraging the model’s self-synthesis capability, MindGYM achieves high data efficiency (e.g., +16% gains on MathVision-Mini with only 400 samples), computational efficiency (reducing both training and inference costs), and robust generalization across tasks.
-Extensive evaluations on seven benchmarks demonstrate superior performance over strong baselines, with notable improvements (+15.77% win rates) in reasoning depth and breadth validated via GPT-based scoring.
-MindGYM underscores the viability of self-challenging for refining VLM capabilities while minimizing human intervention and resource demands.
-Code and data are released to advance multimodal reasoning research.
-**footnotetext: Equal contributions.
+Large foundation models face challenges in acquiring transferable, structured thinking abilities, especially when supervised with rigid templates or crowd-annotated instruction datasets.
+Unlike prior approaches, we focus on a thinking-centric data synthesis paradigm that enables models to evolve through self-generated, cognitively guided data.
+We propose MindGYM, a structured and scalable framework for question synthesis, composed of: (1) Cognitive Thinking Process Injection, which infuses high-level reasoning objectives to shape the model’s synthesis behavior; (2) Seed Single-Hop Question Synthesis, generating atomic questions from diverse semantic types to encourage broader thinking; and (3) Challenging Multi-Hop QA Synthesis, composing more complex multi-hop questions based on QA seeds for deeper reasoning.
+Detailed analysis shows that synthetic data generated by our method achieves 16.7% higher average quality and 67.91% lower quality variance compared to baseline sources, highlighting that both high-quality and self-contained data are essential for effective, thinking-oriented finetuning.
+MindGYM improves performance on six reasoning benchmarks, achieving gains of up to 16% on MathVision using only 400 data samples, and generalizable improvements across different model sizes and architectures.
+MindGYM underscores the viability of self-challenging mechanisms in refining large model capabilities while minimizing human intervention and resource demands.
+Code and data are released to promote data-centric research into self-evolving foundation models driven by their internal reasoning capabilities.
+11footnotetext: Equal contribution.
 1 Introduction
-Large vision-language models (VLMs) have emerged as foundational tools for tasks requiring joint reasoning over visual and textual inputs [4, 30].
-However, enabling robust performance with transferable and efficient reasoning abilities with high-quality data remains challenging [6, 8].
-Manually curated instruction datasets like OK-VQA [25] and ScienceQA [22] are labor-intensive to scale, while self-supervised synthetics methods such as MMInstruct [21] and MMEvol [24] suffer from limited generalization cross modality and tasks, failing to tightly and logically couple visual and textual semantics.
-Meanwhile, reasoning enhancement methods—such as reinforcement learning (RL) [1] or iterative prompting [19]—incur prohibitive computational costs, limiting their practicality.
-To address these challenges, we propose MindGYM, a new framework illustrated in Figure 1 that enhances VLMs via synthetic self-challenging questions with three core stages: (1).
-Seed Single-Hop Question Synthesis: Generates cognitive questions across text-only (e.g., logical deduction) and multimodal contexts (e.g., scientific diagrams), spanning eight semantic topics like causal reasoning and ethical analysis.
-(2).
-Challenging Multi-Hop Question Synthesis: Combines seed questions using cognitive operators—textual strategies (Bridging, Comparison, Temporal) and multimodal reasoning types (Visual-Textual Alignment, Spatial Reasoning, etc.)—to create layered, multi-step problems demanding deeper cognitive engagement.
-(3).
-Thinking-Induced Curriculum Fine-Tuning: A structured training pipeline progresses from scaffolded reasoning (e.g., answer prediction with explicit thought paths) to standalone inference, embedding advanced reasoning capacities into model parameters.
-All these processes are operated on the model to be tuned, via cognitive data self-synthesis and self-challenging training.
-Extensive empirical evaluations across 7 benchmarks demonstrate that MindGYM consistently outperforms prior methods on both text-based and multimodal scenarios, achieving +16% gains on challenging tasks like MathVision-Mini with merely 400 synthesis samples, while reducing computational costs in both training and inference time.
-Notably, MindGYM gains robust generalization across different synthesis sources, and enhanced reasoning depth and breadth via GPT-based scoring analysis over state-of-the-art (SOTA) competitors (+15.77% win rates), enabling the model to handle open-ended tasks requiring multi-modal, high-order cognition.
-Our main contributions can be summarized as follows:
-  • We propose a new cognitive-grounded synthetic data synthesis pipeline, that systematically generates complex questions considering thinking depth, thinking width, and cross-modality.
-  • We show the feasibility of self-challenging evolution for VLMs, with high data-efficiency and progressive training that absorbs explicit human-like reasoning patterns.
-  • We verify MindGYM’s effectiveness, practical efficiency, and contextual adaptability with extensive experiments.
-Our code and data are released at https://github.com/modelscope/data-juicer/tree/MindGYM to foster future progress in vision-language reasoning.
-Figure 1: The proposed MindGYM framework architecture consists of three core stages: ① Seed Single-Hop Question Synthesis, ② Challenging Multi-Hop Question Synthesis, and ③ Thinking-Induced Curriculum Fine-Tuning.
-The process begins with cognitive-aware self-synthesis across text (e.g., mathematical/biological questions) and multimodal (text+image) contexts.
-Seed questions are then combined via text-based strategies (Bridging, Comparison, Temporal) and multimodal reasoning types (Visual-Textual Alignment, Spatial Reasoning, Causal Inference, Contextual Synthesis) to form self-challenging questions.
-Finally, a curriculum-driven fine-tuning phase injects structured cognitive competence into the model through masked-task learning.
-Color-coded regions and arrows illustrate the hierarchical progression from simple reasoning to advanced problem-solving, emphasizing conceptual depth and modal integration.
+Large foundation models have emerged as key assistants for tasks requiring complex understanding across tasks and modalities qwen25vl ; qin2024synergydatamultimodallarge .
+However, enabling robust performance with transferable and efficient thinking abilities remains challenging dj2 ; djsandbox .
+Manually curated instruction datasets like OK-VQA okvqa and ScienceQA lu2022learn are labor-intensive to scale, while self-supervised synthetic methods such as MMInstruct liu2024mminstruct and MMEvol luo2024mmevol suffer from limited generalization across modality and task types, often failing to produce logically consistent or cognitively diverse data.
+Meanwhile, reasoning enhancement methods—such as reinforcement learning (RL) akyurek2023rl4f or iterative prompting kumar2024training —incur prohibitive computational costs, limiting their practicality.
+To tackle these limitations, we introduce MindGYM, a thinking-centric data synthesis framework aimed at enhancing the cognitive capacity of large models.
+Rather than relying on task-specific templates or crowd-sourced samples, our approach embeds structured thinking traits into the synthesis process — enabling models to self-generate data that target their cognitive bottlenecks.
+The proposed framework consists of three key components illustrating in Figure 1:
+  1. 1.
+    Cognitive Thinking Injection: We infuse structured thinking priors — such as breadth (cross-topic linkage), depth (multi-step deduction), and progression (difficulty scaling) — into the prompt design, guiding generation toward cognitively rich and pedagogically effective samples.
+  2. 2.
+    Seed Single-Hop Question Synthesis: We synthesize a set of semantically grounded, single-hop QA primitives that span diverse categories including arithmetic, logic, ethics, and causality.
+These act as composable base units for constructing multi-hop cognitive challenges.
+  3. 3.
+    Challenging Multi-Hop QA Synthesis: By composing seed questions using thinking-centric operators such as Bridging, Comparison, and Temporal, we generate challenging multi-hop questions that demand higher-order inference and cross-domain understanding.
+We validate MindGYM on six reasoning-intensive benchmarks across multiple vision-language models (VLMs) architectures.
+As shown in Table 1, with only 400 synthetic samples, our method yields significant gains: for example, our method boosts performance of Qwen2.5-VL-7B qwen25vl by over 16% on MathVision-Mini.
+These improvements are consistent across models of varying scales and architectures, demonstrating the generality of our approach.
+Further analysis via Data-Juicer dj2 reveals that on Qwen2.5-VL-32B qwen25vl , our synthetic data achieves a +16.7% average quality improvement over baselines and a 67.91% reduction in quality variance.
+This highlights an important insight: not only does cognitively guided synthesis yield higher-quality data, but stability in quality—as measured by low variance—is critical for consistent fine-tuning.
+Moreover, beyond the high-quality textual data synthesis presented in the main page, we also explore MindGYM’s scalability to multimodal settings in Appendix C, demonstrating its effectiveness and potential to both VLMs and large language models (LLMs).
+To summarize, our contributions are three-fold:
+  • We propose MindGYM, a scalable, model-dependent self-synthesis framework for thinking-centric data, which injects structured cognitive priors into both single-hop and multi-hop question generation, encouraging both broad and deep reasoning.
+  • We present an in-depth data analysis, showing that reducing variance is as critical as improving mean quality for stable fine-tuning.
+The proposed method generates cognitively guided data that excels on both fronts.
+  • We demonstrate that self-challenging data evolution significantly enhances reasoning performance across six benchmarks, even under limited supervision, and generalizes well across different models.
+All code and data are released at https://github.com/modelscope/data-juicer/tree/MindGYM/ to shed light on further research and applications.
 2 Related Works
 Instruction Data Construction for VLMs.
-Large vision-language models (VLMs) critically depend on high-quality instruction-tuning datasets to achieve robust performance [imgdiff, 7].
-Early efforts, such as OK-VQA [25] and ScienceQA [22], rely on human-annotated data to ensure accuracy and relevance.
-However, these manual approaches are resource-intensive and difficult to scale.
-To address this limitation, recent works have explored self-instruct paradigms that automate dataset creation.
-While text-centric methods like [39, 34, 42] emphasize diversity and complexity in textual instructions, their extension to multimodal domains remains challenging.
-Multimodal self-instruction frameworks, such as MMInstruct [21] and MMEvol [24], attempt to bridge this gap by iteratively generating image-text pairs.
-Nevertheless, these methods often suffer from modality inconsistency, where generated textual instructions lack tight visual grounding or rely on shallow cross-modal alignments (e.g., generic image descriptions).
-This issue arises due to an over-reliance on LLMs for text generation, which lack explicit mechanisms for visual reasoning.
-Consequently, the resulting datasets fail to challenge VLMs on fine-grained vision-language understanding tasks.
-Reasoning-Aware Question Synthesis.
-Synthetic question generation has shown success in domains like mathematical reasoning [20, 36], where structured problem spaces enable systematic scaling of complexity [32] and skill decomposition [3].
-However, extending these techniques to multimodal domains poses unique challenges due to three key factors: (1) Heterogeneous Semantics: Vision-language reasoning requires handling unstructured inputs with diverse modalities, unlike the formal grammars available in mathematical problems.
-(2) Cross-Modal Synergy: Current methods focus on single-modality skill progression rather than fostering interactions between modalities.
-(3) Open-Ended Challenges: Template-driven approaches, such as Visual Program Induction [12], struggle to generate open-ended reasoning tasks that demand higher-order cognitive skills.
-As a result, existing methods exhibit limited generalizability to complex vision-language tasks, underscoring the need for more flexible and domain-agnostic synthesis frameworks.
-Reasoning-Aware Self-Improvement.
-Enhancing reasoning capabilities in VLMs often incurs significant computational overhead.
-RL-based approaches [1] require extensive policy optimization over large-scale reasoning trajectories [14], making them computationally expensive.
-Iterative refinement techniques [19] further exacerbate costs by introducing multi-stage processing pipelines.
-Even parameter-efficient methods, such as Chain-of-Thought prompting [40], suffer from latency penalties due to lengthy sequential decoding.
-Recent advances in test-time scaling [26] demonstrate promising performance gains but at the expense of increased inference latency.
-These limitations highlight the need for more efficient strategies to improve reasoning without compromising computational feasibility.
-Our Position.
-MindGYM addresses these three critical limitations of prior art: (1) Modality-inconsistency.
-With multiple cognitive relationship enhancement in multi-hop questions, we enforce tight semantic coupling in vision-language concepts.
-(2) Generalization.
-By adopting domain-agnostic prompts, we explicitly incorporate eight vision meta-topics beyond math, and adaptively increase visual grounding and linguistic complexity on the synthesis data.
-(3) Computational efficiency.
-By injecting complex, self-challenging questions directly into the training data, we encourage models through curriculum fine-tuning to perform deep and wide reasoning within a single inference step, eliminating the need for iterative adjustments.
+High-quality instruction data is pivotal for vision-language models (VLMs) to align with human intent jiao2024imgdiffcontrastivedatasynthesis ; kddtutorial .
+While early benchmarks like OK-VQA okvqa and ScienceQA lu2022learn rely on costly human annotation, recent work embraces self-instruction paradigms to scale data generation wang2023self ; taori2023alpaca ; xu2023wizardlm .
+In the multimodal domain, MMInstruct liu2024mminstruct and MMEvol luo2024mmevol extend this idea, but often suffer from modality inconsistency, where visual inputs are weakly grounded to textual outputs.
+These limitations stem from over-reliance on LLMs without structured visual reasoning.
+Our work departs from these trends by embedding cognitively structured priors into prompt design, generating semantically aligned, cross-modal instruction data.
+Thinking-Centric Question Synthesis
+Synthetic data generation has long been a tool for improving model reasoning.
+In language-only domains, question synthesis approaches have emphasized structure, difficulty scaling, and multi-step deduction, particularly in math or logic domains lewkowycz2022solving ; shao2024deepseekmath .
+However, most existing methods either depend on rigid templates or domain-specific structures, which constrains their generalizability.
+Multimodal question synthesis introduces further complexity due to the need for cross-modal reasoning and abstract concept grounding.
+Methods like Visual Program Induction pmlr-v162-duan22c or GRILL parcalabescu2022valse attempt structured generation but often produce shallow or repetitive questions.
+Moreover, many pipelines prioritize syntactic fluency over cognitive depth, failing to elicit the higher-order thinking needed for robust generalization.
+This motivates the need for synthesis frameworks that directly embed cognitive traits—such as reasoning breadth, depth, and progression—into the data itself.
+Self-Improving Reasoning with Limited Supervision
+Self-improvement paradigms for reasoning have emerged to reduce human intervention and labeling costs.
+Reinforcement learning–based methods (e.g., RL4F akyurek2023rl4f ) and iterative self-refinement kumar2024training fine-tune models through repeated exploration and feedback loops but require significant computational resources.
+Lightweight techniques like Chain-of-Thought prompting wei2022chain and Tree-of-Thoughts yao2023tree improve test-time reasoning but often rely on long sequences or search-time strategies, which are inefficient for deployment.
+In contrast, data-centric self-improvement strategies aim to modify the training signal itself rather than the inference process.
+MindGYM fits into this paradigm by synthesizing cognitively aligned, reasoning-targeted samples that directly support robust and efficient learning, without the need for test-time decoding tricks or expensive policy optimization.
 3 The Proposed MindGYM
-In this work, we aim to embed advanced reasoning capabilities into VLMs through synthesized fine-tuning datasets.
-For large models, conventional single-hop question-answer pairs are insufficient to stimulate reasoning in both depth and breadth.
-To address this, we propose a progressive self-synthesis framework, outlined in Figure 1, operating through three interconnected stages:
-(1) Seed Question Synthesis: Starting with a base VLM $\pi$ (e.g., Qwen-2.5-VL [4]), the model self-generates logically related single-hop questions $\{q^{1}_{i},\ldots,q^{k}_{i}\}_{i\in N}$ to explore diverse reasoning methodologies and topics.
-Here, $k$ denotes the number of seed questions per context and $N$ represents the number of synthesized data.
-(2) Self-Challenging Question Synthesis: Using $\{q^{1}_{i},\ldots,q^{k}_{i}\}$ as inputs, the model constructs structured complex questions $\{(Q_{i},A_{i},T_{i})\}_{i\in N}$ via multi-hop combinations.
-Each $Q_{i}$ is formed as $Q_{i}=q^{1}_{i}\oplus\ldots\oplus q^{k}_{i}$, where $\oplus$ represents cognitive operators (e.g., composition, contradiction, abstraction) that create hierarchical reasoning structures.
-The model concurrently generates explicit thinking trajectories $T_{i}$, which documents the rationale behind $Q_{i}$ and $A_{i}$.
-(3) Self-Challenging Cognitive Training: The model is fine-tuned on four structured datasets $\{D^{j}_{\text{train}}\}_{j=1}^{4}$ through a curriculum pipeline.
-Training focuses on predicting masked components across questions $(Q_{i})$, answers $(A_{i})$, and their associated thought paths $(T_{i})$, thereby cultivating deeper and broader reasoning capacity.
-This adversarial questioning approach systematically improves the model’s ability to navigate ambiguous, multi-step reasoning tasks.
-The proposed framework enhances generalization by exposing models to diverse cognitive challenges while injecting interpretable thinking patterns that mirror human-like problem-solving dynamics into the model parameters.
-Detailed design and implementations are provided in subsequent subsections.
+Figure 1: The proposed MindGYM framework incorporates a cognitively guided data synthesis pipeline with four stages: ① Context Generation, ② Single-Hop Question Synthesis, ③ Multi-Hop Composition with Thinking Trace, and ④ Structured Extraction.
+Starting from a meta-topic and guided by a shared cognitive objective, the model iteratively builds background context, atomic reasoning steps, composite questions with interpretable thinking traces, and final structured QA samples for downstream use.
+Color-coded regions and arrows illustrate the hierarchical progression from simple reasoning to advanced problem-solving and emphasizing conceptual depth.
+In this work, we aim to embed advanced thinking capabilities into large models through self-synthesized fine-tuning datasets.
+Prior research has shown that conventional single-hop question–answer pairs are often inadequate to stimulate thinking along both depth and breadth dimensions, especially for tasks that require multi-step deduction or complex reasoning.
+To address this, we introduce a progressive self-synthesis framework, outlined in Figure 1, operating through three interconnected components: (1) Cognitive Thinking Injection, where diverse human-like reasoning strategies are embedded into the synthesis process via structured prompts; (2) Seed Single-Hop Question Synthesis, where the model generates logically grounded, cognitively-aware atomic questions; and (3) Challenging Multi-Hop QA Synthesis, where these atomic units are composed into complex, multi-step questions with explicit reasoning trajectories.
+This framework enhances generalization by exposing models to diverse cognitive challenges while injecting interpretable thinking patterns that mirror human-like problem-solving dynamics into the model parameters.
+Detailed designs and implementations are provided in subsequent subsections.
+3.1 Cognitive Thinking Injection
+Cognitive-Aware Synthesis Templates.
+[…]
+We propose a four-stage cognitive thinking injection framework that incrementally builds structured reasoning traces using a unified generation model $\pi$, guided by $P_{\text{cog}}$ and stage-specific $P_{\text{task}}$, with Stage 1 and Stage 2 described in Section 3.2, and Stage 3 and Stage 4 in Section 3.3:
+  • Stage 1: Background Context Generation.
+Given a meta-topic, the model generates a background passage $P_{\text{context}}$ using $P_{\text{cog}}$ and a context prompt $P_{\text{task}}^{\text{context}}$: $\pi(P_{\text{cog}},P_{\text{task}}^{\text{context}})\rightarrow P_{\text{context}}.$
+  • Stage 2: Single-Hop Question Generation.
+Conditioned on $P_{\text{cog}}$, $P_{\text{context}}$, and $P_{\text{task}}^{\text{single-hop}}$, the model generates $k$ logically independent seed questions: $\pi(P_{\text{cog}},P_{\text{context}},P_{\text{task}}^{\text{single-hop}})\rightarrow\{q_{1},\dots,q_{k}\}.$
+  • Stage 3: Cognitive Composition with Adaptive Types.
+Using the single-hop questions and a multi-hop prompt, the model composes a complex multi-hop question and a corresponding reasoning trace: $\pi(P_{\text{cog}},P_{\text{context}}^{\text{single-hop}},P_{\text{task}}^{\text{multi-hop}})\rightarrow(q_{\text{multi}},\texttt{Thinking}).$
+  • Stage 4: Structured Extraction.
+Given the multi-hop results and a structured prompt, $\pi$ produces the final schema-aligned data: $\pi(P_{\text{cog}},P_{\text{context}}^{\text{multi-hop}},P_{\text{task}}^{\text{structured}})\rightarrow(\texttt{Question},\texttt{Answer},\texttt{Others}).$
+Adaptive Stream of Consciousness.
+To implement $P_{\text{cog}}$, we adapt the Thinking-Claude protocol claude , structuring it as a flexible suite of cognitive operations (e.g., hypothesis formulation and testing, multi-angle interpretation, counterfactual reasoning, and self-verification).
+These thinking elements are modular and reusable, enabling the model to emulate layered, human-like thinking.
+Across stages, $\pi$ is prompted to verbalize and externalize its internal thinking steps, thereby producing interpretable structured traces that reveal how the model navigates complex problem spaces.
+This cognitively guided synthesis approach enhances generalization by exposing the model to diverse reasoning challenges while injecting structured and interpretable thinking into training data.
+The resulting dataset contains rich cognitive signals that are later leveraged during fine-tuning (see Section 3.4).
+Empirical evidence for the efficacy of $P_{\text{cog}}$ is presented in Section 4.4.
 3.2 Seed Single-Hop Question Synthesis
-3.1.1 Diverse Thinking Methodologies
-Cognitive-Aware Synthesis Templates.
-We adopt a synthesis template that injects thinking elements, consisting of three parts: cognitive setting, background conditions, and synthesis goals: $P=\{P_{cog},P_{context},P_{task}\}$.
-$P_{context}$ and $P_{task}$ have different instantiations in different synthesis scenarios (single-hop, multi-hop QA, text, multimodal) in our method  [^1] The full prompt is presented in the Appendix B., while $P_{cog}$ remains unchanged to encourage the model to consistently produce responses that are strongly related to thinking.
-We first describe the single-hop question scenario, with other tasks detailed later.
-Based on the given $P_{context}$, $P_{task}$ requires $\pi$ to generate $k$ logically connected seed single-hop questions $q^{1}_{i},...q^{k}_{i}$, at the same time, the model will be explicitly noted that these seed questions will be used later as more complex multi-hop combinations (the specific combination types will be detailed in later Sec.3.2).
-Adaptive Stream of Consciousness.
-We draw inspiration from a thinking protocol named Thinking-Claude [44] and adapt it to shape $P_{cog}$ as a collection of many diverse thinking processes, (e.g., making and challenging hypotheses, exploring and balancing different perspectives).
-This makes the model mimic the human thinking process in a layered, context-aware style, aiming for a broader coverage in thinking methodologies.
-Given different $P_{context}$ and $P_{task}$, $\pi$ dynamically adjusts thinking dimensions, such as problem depth, interpretation angles, and validation rigor, and explicitly expose models’ “minds” dealing with complex scenarios.
-This adaptive exploration not only encourages the model to incorporate diverse reasoning paths for generalizing to reasoning tasks, but also injects structured variability into training data, countering the memorization biases in conventional pipelines.
-Empirical validation of $P_{cog}$ is presented in Section 4.3.
-3.1.2 Diverse Thinking Objects
-Meta-Topics.
-We consider both text-only and text-image QA synthesis.
-Here we first introduce the $P_{context}$ for the text-only case.
-Building on the cognitive architecture theory [28] and the multi-compartmental knowledge representation [33], we establish eight orthogonal meta-topics spanning four foundational reasoning dimensions: quantitative, causal, temporal, and social-ethical.
-Unlike prior works focused on limited domain combinations [2], our approach emphasizes broad cognitive coverage:
-  • Mathematical Reasoning: Solving quantitative problems using numbers and formulas.
-  • Scientific Knowledge: Understanding natural laws and scientific principles.
-  • Logical Deduction: Forming conclusions through logical progression from premises.
-  • Technical Procedures: Following step-by-step instructions for practical tasks.
-  • Historical Events: Analyzing past events and their consequences.
-  • Ethical Considerations: Evaluating decisions through moral frameworks.
-  • Economic Principles: Studying resource allocation and valuation.
-  • Psychological Insights: Exploring human behavior and cognition.
-These eight topics are selected for two core reasons: (1) alignment with widely studied reasoning domains (e.g., mathematics, logic) [17], and (2) extension to synergize with VLMs’ capabilities and potential [30], balancing foundational reasoning with subsequent self-challenges.
-By polling one of these meta-topics and their descriptions, we require $\pi$ to self-generate background document as $P_{context}$, then self-generate $k=5$ seed single-hop questions per inference.
-Empirical results in Section 4.3 demonstrate that this topic-balanced sampling strategy enhances thinking breadth and significantly improves end-to-end model performance compared to unbalanced alternatives.
-This process allows $\pi$ to explore maximal topic diversity, leveraging its internal knowledge to identify domains and skills requiring advanced reasoning.
-Anchoring Image Sources.
-For the $P_{context}$ component in text-image QA synthesis, we use existing images rather than generating images from scratch (as in the text-only case).
-This is because the existing SOTA VLMs are mostly only image-to-text generation (i.e., image understanding) and fail to do text-to-image generation at the same time.
-This decision arises because current SOTA VLMs predominantly specialize in image-to-text generation (i.e., image understanding) and lack dual capabilities for text-to-image generation constrained by their architectures.
-Without anchored visual references, the model risks hallucinatory reasoning about image elements (e.g., referencing “growth curves” might lead to incorrect inferences about unrelated scenarios, such as spatial-based reasoning tasks).
-Specifically, we randomly sample image-text pairs from scientific knowledge and general-domain QA datasets – ScienceQA [22] and OKVQA [25] – to serve as background context ($P_{\text{context}}$) for $\pi$.
-We emphasize alignment between visual and textual concepts within the generation instruction $P_{\text{task}}$, prompting $\pi$ to self-generate $k=5$ seed questions per inference.
-This enhances harmonized interactions between image and text modalities during synthesis.
+Cognitively Grounded Meta-Topics.
+To ensure reasoning diversity and depth, we construct eight orthogonal meta-topics derived from established cognitive theories, including unified cognitive architectures newell1994unified and dual-representation models sun2001duality .
+These meta-topics span four core reasoning dimensions: quantitative, causal, temporal, and social-ethical, offering a comprehensive blueprint for cognitively rich question generation.
+Detailed meta-topics are provided in Appendix B.2.
+Stage 1: Background Context Generation.
+Given a sampled meta-topic and its brief description, the model is prompted with a cognitive grounding prompt $P_{\text{cog}}$ and a context-specific instruction $P_{\text{task}}^{\text{context}}$ to self-generate a short but semantically rich background passage $P_{\text{context}}$.
+This passage serves as a reference for subsequent single-hop question generation.
+Stage 2: Single-Hop Question Generation.
+Conditioned on the generated background passage $P_{\text{context}}$, the cognitive intent $P_{\text{cog}}$, and a task-specific instruction $P_{\text{task}}^{\text{single-hop}}$, the model generates a small batch (up to $k=5$) of logically grounded, atomic single-hop questions $\{q_{1},q_{2},\dots,q_{k}\}$.
+These questions are designed to probe individual reasoning steps based on the passage and are explicitly treated as intermediate components for later multi-hop composition in Section 3.3.
 Reject Sampling for Diversity.
@@
 We use an empirically determined cosine similarity threshold to evaluate the overlap between newly generated and existing samples.
-If similarity exceeds this threshold, the system initiates a regeneration cycle to eliminate semantically redundant content, continuing until the seed question pool reaches the target size $N$.
-This mechanism widens the breadth of reasoning in synthetic data, better preparing it for diverse downstream scenarios.
-Unlike prior methods concentrating on domain-specific reasoning (e.g., mathematical tasks), MindGYM expands reasoning scope across three dimensions: cognitive-aware thinking methodologies, image-text interactions, and semantic redundancy filtering.
-In Section 4.2, we empirically validate MindGYM ’s robustness and effectiveness across varied synthesis sources—including model’s internal knowledge, multilingual contexts, and anchored visual content.
+If similarity exceeds the threshold, the system initiates a regeneration cycle to eliminate semantically redundant content, continuing until the seed question pool reaches the target size $N$.
+This reject sampling mechanism widens the breadth of reasoning in synthetic data.
+Unlike prior works that narrowly focus on limited task domains austin2021program ; huang2023towards (e.g., arithmetic, logic puzzles), MindGYM emphasizes (1) internalized background grounding, (2) cognitively structured topic guidance, and (3) semantic diversity enforcement.
+In Section 4.4, we empirically validate MindGYM robustness and effectiveness across varied synthesis sources, including model’s internal knowledge and multimodal contexts.
 3.3 Challenging Multi-Hop QA Synthesis
-3.2.1 Cognitive Combination Principles
-Synthesis Templates for In-depth Thinking.
-Recall that in Section 3.1, we generate $k$ single-hop seed questions and proactively inform $\pi$ that these questions will be used for future cognitive combinations.
-Here, we explicitly specify some cognitive combination relationships into the synthesis requirement $P_{task}$ for multi-hop QAs, and let $\pi$ to self-challenging by figuring out how to combine the generated $\{q^{1}_{i},...q^{k}_{i}\}_{i\in N}$ into more complex questions $\{Q_{i}\}_{i\in N}$ conditioned on aforementioned $P_{cog}$ and $P_{context}$, such that the results is logically right, cognitively considered and match the most suitable relationships among the given ones.
-Building on the single-hop seed questions generated in Section 3.1, we advance to multi-hop question synthesis by explicitly defining cognitive combination relationships within the task specification $P_{\text{task}}$.
-During this phase, the model $\pi$ is prompted to synthesize more complex questions $\{Q_{i}\}_{i\in N}$ by logically combining the pre-generated seed questions $\{q^{j}_{i}\}_{i=1,...,N;j=1,...,k}$ under constraints from $P_{\text{cog}}$ (cognitive framework) and $P_{\text{context}}$ (factual reference).
-This process aims to make $\{Q_{i}\}_{i\in N}$ logically valid, cognitively grounded, and optimally match the most relevant relationships from our predefined taxonomy.
-Textual Cognitive Combination Types.
-For $P_{\text{context}}$ containing only textual content, we define three cognitive combination forms after initial seed generation:
-  • Comparison: Requires contrasting two or more concepts or scenarios.
-  • Bridging: Connects disparate concepts through intermediate reasoning steps.
-  • Temporal: Involves analyzing events across chronological contexts.
-Multimodal Cognitive Combination Types.
-For $P_{\text{context}}$ containing both text and images, we introduce four multimodal combination types:
-  • Visual-Textual Alignment: Ensures consistency between visual data (e.g., diagrams, charts) and textual descriptions.
-  • Spatial Reasoning: Addresses spatial relationships or geometric arrangements depicted in visual content.
-  • Causal Inference: Requires deducing cause-effect relationships from combined textual and visual inputs.
-  • Contextual Synthesis: Demands integration of information across modalities (e.g., text, images, graphs) to form unified conclusions.
-3.2.2 Cognitive Manifestation
-Structured Cognitive Processes & Reasoning Units.
-Beyond specifying cognitive combination relationships in the task requirement $P_{\text{task}}$, we explicitly require $\pi$ to articulate its intermediate thought process $T_{i}$ leading to the final question $Q_{i}$.
-This support $T_{i}$ links the seed questions $\{q^{j}_{i}\}_{i\in N}$ into a logically coherent reasoning pathway, explicitly connecting steps to ensure interpretability and contextual consistency.
-To balance cognitive context relevance and inference efficiency, $\pi$ is prompted to jointly generate $Q_{i}$, $A_{i}$, and $T_{i}$ in one inference step.
-However, raw responses may include fragmented or overly lengthy content, so we further prompt $\pi$ to segment outputs into distinct components: formal question-answer pairs $(Q_{i},A_{i})$ and their supporting rationale $T_{i}$.
-This results in modular cognitive units of varying granularity, enabling more flexible refinement of $\pi$’s reasoning abilities in subsequent training (more details are in later Section 3.3).
-Figure 2: An end-to-end example of MindGYM.
-For multimodal data, we generate five seed questions first, and make the model self-challenging itself via synthesizing multi-hop questions and multi-hop answers while preserving its internal thinking process.
-Cognitive Information Processing Dynamics.
-Note that the synthesis framework explicitly involves $\pi$ processing both novel and existing knowledge during inference.
-By reintroducing context elements from $P_{\text{context}}$, accessing seed questions $\{q^{j}_{i}\}_{i\in N}$, and addressing advanced combination requirements in $P_{\text{task}}$, the model is incentivized to: (1).
-Explore diverse reasoning patterns embedded in $P_{\text{cog}}$, (2).
-Re-examine context details with renewed emphasis, and (3).
-Construct multi-hop answers using enriched relational constraints.
-This process yields more focused, nuanced questions $Q_{i}$ while eliminating reasoning shortcuts, forcing $\pi$ to engage in deeper inferential analysis.
-Figure 2 illustrates an example in which the original $P_{\text{context}}$ (including image and textual inputs) is first used to synthesize seed questions (e.g., “What ingredients did Kenji use?” or “What vitamins are in broccoli?”), as detailed in Section 3.1).
-These are then combined via the multi-hop synthesis protocol (outlined in Section 3.2) into more challenging questions that require cross-contextual reasoning (e.g., “How do Kenji’s lunch ingredients compare with Tracy’s in terms of vitamin content?” paired with bridging and temporal relationships).
-Critically, whereas the original QA focused on limited details (e.g., Kenji’s tomatoes and Tracy’s broccoli), our synthesized questions expand reasoning depth by incorporating broader contextual elements: sandwich compositions, citrus fruits, root vegetables, etc.
-The resulting multi-hop QAs demonstrate richer interconnections, broader scope, and higher cognitive demands on the solver.
-3.3 Thinking-Induced Curriculum Fine-Tuning
-Training Objectives.
-Given the self-synthesized dataset containing triples $\{(Q_{i},A_{i},T_{i})\}_{i\in N}$, we propose a “MindGym”-style training flow to progressively integrate cognitive guidance into the model $\pi$ through mixed masking strategies.
-This approach incrementally challenges $\pi$ by varying the exposure to question-answer pairs ($Q_{i}$, $A_{i}$), their supporting rationales ($T_{i}$), and masked input-output configurations.
-For recap, $Q_{i}$ represents the complex multi-hop question, $A_{i}$ denotes the reference answer with structured reasoning, and $T_{i}$ records the detailed step-by-step cognitive trajectory.
-We construct four training datasets with distinct masking objectives:
-$\mathcal{D}^{1}_{\text{train}}=\{(Q_{i},T_{i}\rightarrow A_{i})\}$ employs the masked input structure $\langle Q_{i},T_{i},[\mathbf{\text{MASK}}]\rangle$.
-The model predicts $A_{i}$ using $Q_{i}$ and explicit thinking guidance $T_{i}$, and the loss is defined as:
-$\mathcal{L}_{1}=-\mathbb{E}_{(Q,T,A)\sim\mathcal{D}^{1}_{\text{train}}}\log\pi_{\theta}(A|Q,T).$ (1)
-$\mathcal{D}^{2}_{\text{train}}=\{(Q_{i},A_{i}\rightarrow T_{i})\}$ uses the structure $\langle Q_{i},[\mathbf{\text{MASK}}],A_{i}\rangle$.
-The objective is to recover the hidden thinking path $T_{i}$, with loss:
-$\mathcal{L}_{2}=-\mathbb{E}_{(Q,A,T)\sim\mathcal{D}^{2}_{\text{train}}}\log\pi_{\theta}(T|Q,A).$ (2)
-$\mathcal{D}^{3}_{\text{train}}=\{Q_{i}\rightarrow(A_{i},T_{i})\}$ follows $\langle Q_{i},[\mathbf{\text{MASK}}]\rangle$, requiring simultaneous generation of $A_{i}$ and $T_{i}$:
-$\mathcal{L}_{3}=-\mathbb{E}_{(Q,A,T)\sim\mathcal{D}^{3}_{\text{train}}}\log\pi_{\theta}(A,T|Q).$ (3)
-$\mathcal{D}^{4}_{\text{train}}=\{Q_{i}\rightarrow A_{i}\}$ uses the same masking strategy as $\mathcal{D}_{train}^{3}$, with input $\langle Q_{i},[\mathbf{\text{MASK}}]\rangle$.
-This mimics standard supervised fine-tuning practice:
-$\mathcal{L}_{4}=-\mathbb{E}_{(Q,A)\sim\mathcal{D}^{4}_{\text{train}}}\log\pi_{\theta}(A|Q).$ (4)
-The curriculum progresses through these datasets sequentially: $\mathcal{D}^{1}\rightarrow\mathcal{D}^{2}\rightarrow\mathcal{D}^{3}\rightarrow\mathcal{D}^{4}$.
-This mirrors a guided learning pathway where $\pi$ first reproduces answers with external cognitive scaffolding ($\mathcal{D}^{1}$), then deduces thinking paths without full guidance ($\mathcal{D}^{2}$), gradually internalizes end-to-end reasoning ($\mathcal{D}^{3}$), and finally masters standalone Q-to-A mappings ($\mathcal{D}^{4}$).
-Empirical validation in Sec.
-4.3 shows this strategy outperforms non-curriculum mixed training.
-Discussion on Efficiency and Cost.
-The MindGYM framework enhances $\pi$’s reasoning capabilities through $N\times 4$ synthesis inferences across stages ($P_{\text{context}}$, seed questions, $Q_{i}$, and $\mathcal{D}_{\text{train}}$).
-Despite this incremental setup, inference efficiency is maintained through amortized reasoning costs during training.
-A single inference suffices to achieve high model performance, as shown empirically in Section 4.3.
-Unlike test-time scaling approaches [38, 18], we move cognitive enhancement costs to the training phase.
-Compared to data-intensive methods (e.g., reinforcement learning) requiring extensive trial-and-error [19], our method maximizes cognitive depth leveraging structured synthetic data without resource-heavy exploration phases.
-Datasets Multimodal Eval Text-Only Eval Avg
-MMStar MathVista MathVision MM-Avg GSM8K MATH GPQA Text-Avg
-Baslines Raw 64.00 69.30 24.67 52.66 83.62 67.60 31.83 61.02 56.84
-LIMO (817) [43] 63.93 70.20 25.9 53.37 84.08 67.80 30.58 60.82 57.10
-Open-O1 (77k) [35] 58.93 61.40 17.43 45.92 77.94 59.00 28.19 55.04 50.48
-Open-O1 (400) [35] 63.87 69.70 23.66 53.08 84.31 69.00 29.7 61.00 57.04
-MMEvol-SciQA (106) [24] 63.93 69.50 23.36 52.26 83.40 65.60 31.83 60.28 56.27
-MMEvol-DVQA (4k) [24] 62.93 67.40 24.01 51.45 83.85 65.80 33.08 60.91 56.18
-Text MindGYM (400) 64.33 70.30 28.62 54.42 84.08 68.40 31.33 61.27 57.84
-Image MindGYM-OKVQA (400) 63.60 69.20 28.29 53.70 84.00 68.20 31.08 61.09 57.40
-MindGYM-SciQA (400) 64.20 69.80 25.33 53.11 83.93 69.00 31.45 61.46 57.29
-Table 1: Performance of Qwen2.5-VL-7B on various evaluation benchmarks. The number following the dataset name indicates its size in terms of the number of samples used for training.
+Stage 3: Cognitive Composition with Adaptive Types.
+Given the seed questions $\{q^{j}_{i}\}_{j=1}^{k}$ from Section 3.2, we treat them as updated context $P_{\text{context}}^{\text{single}}$ for this phase.
+The model $\pi$ is now prompted with the triplet $(P_{\text{cog}},P_{\text{context}}^{\text{single}},P_{\text{task}}^{\text{multi-hop}})$ to synthesize a more complex, cognitively grounded question $Q_{i}$.
+The task instruction $P_{\text{task}}^{\text{multi-hop}}$ specifies an abstract reasoning requirement (e.g., comparing entities, bridging concepts, or analyzing temporal relations), which guides the model to self-compose from prior seeds.
+Details of composition types are provided in Appendix B.
+Stage 4: Structured Extraction.
+To support modular training and interpretation, we encourage $\pi$ to format its output in three clearly separated blocks: the final multi-hop question $Q_{i}$, its answer $A_{i}$, and the structured rationale $T_{i}$.
+These thinking traces serve as valuable training signals for reasoning-intensive QA tasks, which is used for training in Section 3.4).
+3.4 Usage and Discussion
+Training with the Dataset.
+In this section, we explore how to effectively utilize our cognitively annotated dataset $\{(Q_{i},A_{i},T_{i})\}_{i\in N}$ by organizing the training process into a gradually evolving pathway that transitions from explicit external guidance ($T_{i}$) to self-contained reasoning.
+The process begins with guided answering, where the model receives both the question and the reasoning trace ($Q_{i}$, $T_{i}\rightarrow A_{i}$) to encourage answer generation aligned with structured thought.
+It then advances to reason reconstruction, where the model is given the question and answer and learns to infer the underlying rationale ($Q_{i}$, $A_{i}\rightarrow T_{i}$), reinforcing logical coherence.
+Next comes paired reasoning, which requires the model to jointly generate both the answer and rationale from the question alone ($Q_{i}\rightarrow A_{i}$, $T_{i}$), simulating independent reasoning under minimal supervision.
+Finally, in the autonomous solving phase, the model learns to directly produce the answer from the question ($Q_{i}\rightarrow A_{i}$), fully internalizing the reasoning process.
+This learning pathway reflects the natural progression of human cognitive development—starting from external support and moving toward self-directed problem solving.
+As demonstrated in Section 4.4, this structured progression leads to significantly better performance compared to training with uniform supervision.
+Extensibility of the Dataset.
+In addition to our primary focus on cognitively annotated textual data, we conduct preliminary experiments on extending our synthesis framework to the multimodal setting.
+As described in Appendix C, we leverage VLMs to generate multi-hop QA samples grounded in visual context.
+Specifically, we use OK-VQA okvqa and ScienceQA lu2022learn as anchor image sources to guide the synthesis process.
+The generated results summarized in Appendix C.3 show encouraging quality and reasoning depth, suggesting the potential of our framework beyond the text-only domain.
+Nonetheless, multimodal synthesis with VLMs remains an open challenge.
+Current reliance on static image datasets as anchors limits the diversity and scalability of visual reasoning.
+Looking forward, we plan to (1) incorporate richer and more diverse image datasets to better stimulate cognitive engagement, and (2) further explore fully generative pipelines where VLMs synthesize both the image and the associated reasoning task, enabling models to improve through self-generated multimodal experiences—a key step toward the self-evolution of models.
+Table 1: Performance of different models on various evaluation benchmarks. The number following the dataset name indicates its size in terms of the number of samples used for training.
+Text Eval Multimodal Eval
+Models Dataset (# of Samples) GSM8K Math GPQA MMStar MathVista MathVision Overall
+[…]
+Openo1-sft (400) openo1 84.31 69.00 29.70 63.87 69.70 23.66 57.04 (+0.20)
+Openo1-sft (4k) openo1 77.94 59.00 28.19 58.93 61.40 17.43 50.48 (-6.36)
+[…]
+MMEvol-SciQA (106) luo2024mmevol 83.40 65.60 31.83 63.93 69.50 23.36 56.27 (-0.57)
+MMEvol-DvQA (4k) luo2024mmevol 83.85 65.80 33.08 62.93 67.40 24.01 56.18 (-0.66)
+[…]
+[…]
+[…]
+[…]
+[…]
+MMEvol-SciQA (106) luo2024mmevol 95.60 81.00 42.42 69.00 73.80 36.84 66.44 (-1.14)
+[…]
+[…]
+raw 86.50 76.30 33.84 69.00 73.20 33.22 62.46
+Openo1-sft (400) openo1 89.39 77.20 43.43 69.00 73.60 32.57 64.20 (+1.74)
+[…]
+[…]
+[…]
+MMEvol-DvQA (4k) luo2024mmevol 88.55 74.60 40.40 68.47 71.60 32.24 62.65 (+0.19)
+[…]
+raw 89.16 72.60 47.47 72.40 72.80 36.51 65.16
+[…]
+Openo1-sft (4k) openo1 89.46 79.20 44.95 72.47 72.90 35.53 65.75 (+0.59)
+LIMO (817) ye2025limo 89.16 72.60 47.47 72.40 72.80 36.51 65.16 (+0.00)
+[…]
+MMEvol-DvQA (4k) luo2024mmevol 89.92 74.00 46.46 70.93 71.70 30.92 63.99 (-1.17)
+[…]
 4 Experiments
-To comprehensively evaluate the performance of our proposed synthetic data method, we conduct extensive experimental validation on multiple representative text and multimodal evaluation sets, comparing it with several competitive data synthesis methods and datasets.
+To comprehensively evaluate the performance of our proposed synthetic data method, we conduct extensive experimental validation on multiple representative evaluation sets.
 Full descriptions of datasets, benchmarks, implementation details, and baselines are provided in Appendix A.
 4.1 Experimental Settings
+Models & Implementation.
+The data synthesis in MindGYM leverages VLMs’ inherent capabilities for long-context understanding and reasoning.
+We adopted four vision language models to verify the generality of our data synthesis method.
+These models are selected to cover both scaling within a single model family (Qwen2.5-VL 7B and 32B) and variation across different models (Qwen and InternVL series).
+For all experiments in the main text, we focus on text-only data synthesis, where only the LLM-layers are updated during training, while all vision-related layers remain frozen.
 Datasets.
-We conduct multi-modality data synthesis across textual and image contexts.
-For textual synthesis, data generation occurs exclusively through self-contained knowledge extraction from the base model, eliminating external data dependencies.
-Multimodal synthesis utilizes established scientific and reasoning corpora as visual primitives, with visual grounding from ScienceQA [22] and OK-VQA [25].
+We focus exclusively on synthetic textual data generation in the main body of our work, relying solely on self-contained knowledge extraction from the base model without external data sources.
+To verify the robustness and generality of our synthesis method, we employ the aforementioned models to perform self-consistent data generation.
+Each model independently synthesizes 400 Chinese samples.
+The choice of Chinese is driven by its linguistic complexity and higher information density, making the synthesis task more challenging.
 Benchmarks.
-To systematically assess MindGYM on VLMs across reasoning modalities, we incorporate:
-  • Multimodal evaluation sets include MMStar [9], MathVista-Mini [23], and MathVision-Mini [37], designed to comprehensively assess the reasoning abilities in the multimodal scenario.
-  • Text-based evaluation sets comprise GSM8K [10], MATH [15], and GPQA [31], focused on evaluating reasoning and mathematical capabilities.
-The MM-Avg and Text-Avg metrics serve as unified criteria for assessing multimodal and text-based reasoning capabilities respectively, whereas Avg quantifies comprehensive performance spanning both unimodal and multimodal reasoning domains.
-Models & Implementation.
-The data synthesis of MindGYM leverages VLMs’ inherent long-context understanding and reasoning capabilities.
-We implement Qwen2.5-VL-7B-Instruct [4] as our base VLM since alternative models (e.g., InternVL2.5-7B) exhibit inferior generation quality under identical configurations (detailed in Appendix D).
-To maintain linguistic consistency with Qwen2.5-VL-7B-Instruct and consider the information density of reasoning data, we utilize Chinese for data synthesis.
-All experiments follow consistent protocols: during text-only training, only LLM-layers update parameters, while image-text joint training additionally activates Aligner-layers, keeping ViT-layers permanently frozen.
+To systematically assess MindGYM on VLMs across reasoning modalities, we incorporate: (1) Text-based evaluation sets comprise GSM8K gsm8k , MATH math , and GPQA gpqa ; (2) Multimodal evaluation sets include MMStar MMStar , MathVista-Mini mathvista , and MathVision-Mini mathvision .
+The Overall quantifies comprehensive performance spanning both text and multimodal domains.
 Baselines.
 As MindGYM generates self-contained adversarial training data without external sources, direct comparison with conventional data synthesis methods proves infeasible.
-We therefore select SOTA reasoning-oriented dataset works from two categories: (1) Multimodal reasoning using MMEvol’s [24] core subsets ScienceQA and DVQA, and (2) Text-based reasoning with LIMO [43], 817 curated logical chains, and Open-O1 [35], an SFT dataset for CoT activation as fundamental text reasoning baselines.
+We therefore select SOTA reasoning-oriented dataset works from two categories: (1) Text-based reasoning with LIMO ye2025limo , 817 curated logical chains, and Open-O1 openo1 , an SFT dataset for CoT activation as fundamental text reasoning baselines, and (2) Multimodal reasoning using MMEvol’s luo2024mmevol core subsets ScienceQA and DVQA.
+All baseline datasets are used in accordance with their original supervised fine-tuning setting.
+For fair comparison, we apply identical training protocols and hyperparameter configurations across all methods.
 4.2 Main Results
 To validate the efficacy of our methodology, we systematically evaluate the model across multiple benchmarks as documented in Table 1, with key insights summarized:
-Universal Superiority.
-MindGYM demonstrates substantial performance advantages through comprehensive average metric evaluation in MindGYM-Text with a 54.42 MM-Avg-score and 57.84 Avg-score, and notably achieving 28.62 on MathVision with a 16% improvement over the Raw model.
-The proposed method significantly outperforms all baselines where even the strongest competitor LIMO trails by 1.05 points in MM-Avg.
-An interesting phenomenon emerges among baselines that Open-O1 (77K) suffers severe degradation with 6.74-point MM-Avg decline and 29.3% MathVision collapse despite employing a large number of training samples, exposing the vulnerability of conventional scaling approaches.
-Cross-modal Transfer.
-MindGYM enables effective cross-modal reasoning capability sharing through bidirectional knowledge transfer.
-Remarkably, the text-specialized MindGYM-Text establishes new SOTA performance on MM-Avg, with the score of 54.42 that surpasses all baselines.
-Conversely, the multimodal-enhanced MindGYM-SciQA demonstrates unexpected supremacy in textual reasoning tasks, achieving an optimal Text-Avg score of 61.46.
-This bidirectional enhancement confirms the transferability of reasoning capabilities between textual and multimodal representations.
-Generalizability & Robustness.
-Consistent improvements emerge across diverse data sources.
-Our approach attains sub-optimal average performance when synthesizing data from both SciQA, showing an improvement of 0.45 over raw data, and OKVQA, with an improvement of 0.56 over raw data, demonstrating adaptability to various domain-specific reasoning tasks.
-This indicates that our method effectively enhances generalizability and robustness across diverse reasoning challenges.
+Superiority across Different Tasks.
+MindGYM demonstrates substantial performance advantages through comprehensive average metric evaluation.
+Notably, on the Qwen2.5-VL-7B, MindGYM achieves the Overall of 57.84, outperforming all competing baselines.
+It achieves the highest MathVision score (28.62), exceeding the Raw model by 16% and the strongest baseline LIMO by 2.72 points.
+An interesting phenomenon emerges among baselines that Open-O1 (4K) suffers severe degradation with 29.3% MathVision collapse despite employing a large number of training samples, exposing the vulnerability of conventional scaling approaches.
+Superiority on Different Models.
+MindGYM consistently demonstrates strong robustness across both diverse models and varying parameter scales.
+On the Qwen2.5-VL series, MindGYM improves the Overall performance from 57.84 to 64.33 on the 7B model, a relative gain of +0.74%–7.36%, and from 68.06 to 68.06 on the 32B model, reflecting a more modest but stable +0.48%–3.31% improvement.
+Similarly, for InternVL3, the Overall score rises from 64.95 to 64.95 on the 14B model (+0.12%–2.30%), and from 66.23 to 66.23 on the 38B model (+0.40%–2.24%).
+These gains remain consistently positive across all four backbones, even as baseline capabilities increase with scale.
+Notably, the relative improvements are more substantial at smaller model sizes (e.g., Qwen2.5-VL-7B), suggesting MindGYM is particularly effective in lower-resource or less capable models.
+This trend underscores the method’s architecture-agnostic and scaling-friendly design, making it highly transferable across a wide range of multimodal foundation models.
 Data Efficiency.
-MindGYM actives model potential using only 400 synthesized samples, achieving a 28.62 MathVision score that surpasses baselines trained with 10$\times$-100$\times$ more data.
-This contrasts sharply with Open-O1’s performance collapse despite massive data scaling, indicating that MindGYM significantly enhances the model’s reasoning capabilities while also ensuring that the model’s performance is not detrimentally affected.
-These findings effectively validate that our approach addresses the issues of modality inconsistency, generalization, and computational efficiency mentioned in Section 2.
+Despite using only 400 synthetic training samples, MindGYM achieves performance that not only matches but exceeds baselines trained with 10$\times$ more data.
+This includes outperforming Open-O1 (4k) in both textual and multimodal tasks.
+On MathVision, MindGYM reaches 28.62 while Open-O1 (4k) collapses to 17.43.
+This stark contrast underscores the high quality and utility of the data synthesized by MindGYM.
+Cross-modal Enhancement.
+Despite containing only 400 synthetic examples in pure-text format, MindGYM significantly enhances both textual and cross-modal reasoning capabilities.
+As shown in Table 1, our method improves the Overall performance not only on language-only tasks, but also on multimodal benchmarks (e.g., from 62.46 to 64.95 on InternVL3-14B and 65.16 to 66.23 on InternVL3-38B).
+These consistent improvements are also observed in vision-intensive benchmarks like MathVista and MathVision, even though no vision data was used in fine-tuning.
+This demonstrates that our synthetic dataset effectively strengthens the model’s intrinsic thinking capability, enabling better generalization across modalities.
+Such cross-modal enhancement, achieved through text-only supervision, highlights the thinking-centered nature of our data and its ability to propagate improvements to both language and vision domains.
+Summary.
+These results collectively demonstrate that our method effectively tackles three key challenges in multimodal instruction tuning: data inefficiency, modality transferability, and scalability across different models and sizes.
+First, MindGYM achieves consistent and significant gains across all evaluation metrics with only 400 synthetic Chinese text-only samples, outperforming baselines trained with up to 10× more data, highlighting its exceptional data efficiency.
+Second, the improvements span both textual and vision-intensive benchmarks, even without visual supervision, indicating that our thinking-centric text data successfully enhances intrinsic thinking that transfers across modalities.
+Third, our method remains robust and effective across a range of backbones (Qwen2.5-VL and InternVL3) and model sizes (7B–38B), making it highly scalable and architecture-agnostic.
+These findings validate that our approach addresses the limitations of prior methods in terms of modality inconsistency, generalization, and computational efficiency (as outlined in Section 2).
+4.3 Data analysis
+Table 2: Data-Juicer analysis results across different models.
+Model Dataset
+quality
+-mean
+$\uparrow$
+quality
+-std
+$\downarrow$
+action $\uparrow$ dependency $\uparrow$ token length
+baseline Openo1-sft (400) openo1 0.96 0.091 8.05 2.00 77 274
+Openo1-sft (4k) openo1 0.70 0.10 8.78 2.02 83 284
+LIMO (817) ye2025limo 0.87 0.10 7.71 2.02 101 322
+MMEvol-SciQA (106) luo2024mmevol 0.91 0.082 2.31 1.85 13.73 67
+MMEvol-DvQA (4k) luo2024mmevol 0.76 0.11 2.66 1.95 14.82 69
+Qwen2.5-VL-7B MindGYM-Text (ours) 0.93 0.055 10.34 2.17 147 110
+Qwen2.5-VL-32B MindGYM-Text (ours) 0.98 0.031 10.94 2.25 168 134
+InternVL-8B MindGYM-Text (ours) 0.97 0.044 8.64 2.16 233 195
+InternVL-38B MindGYM-Text (ours) 0.96 0.039 7.66 2.07 118 92
+To comprehensively assess the quality of our MindGYM synthetic dataset, we adopt Data-Juicerdj2  [^2] https://github.com/modelscope/data-juicer, a modular data-centric analysis toolkit that provides LLM-guided operators for probing data across various dimensions.
+Specifically, we compare MindGYM against several baselines (e.g., LIMO, Open-O1, MMEvol) using five diagnostic filters: (1) quality, which estimates the overall textual quality based on LLM judgment; (2) action, counting the number of action verbs as a proxy for instruction richness; (3) dependency, which penalizes the presence of non-independent noun phrases based on syntactic trees; (4) token, evaluating token-length consistency; and (5) length, which considers raw text length variability.
+Table 2 summarizes the full results across multiple model series and checkpoints.
+Our key findings are elaborated below.
+Table 3: The results of win rates in terms of relative improvements on MindGYM over baselines, using GPT-4 as a scorer.
+Datasets Depth Breadth Avg
+Raw 10.2% 19.4% 14.8% $\uparrow$
+LIMO 1.63% 23.8% 12.7% $\uparrow$
+Open-O1 (400) 1.88% 23.7% 12.8% $\uparrow$
+MMEvol-DVQA 8.41% 37.0% 22.7% $\uparrow$
+Overall 5.53% $\uparrow$ 26.0% $\uparrow$ 15.8% $\uparrow$
+Higher Quality.
+We observe a strong positive correlation between the average quality score (as estimated by Data-Juicerdj2 ) and downstream finetuning performance.
+In the baseline group, Open-O1(400) and MMEvol-SciQA achieve high quality scores (0.96 and 0.91, respectively), and correspondingly rank first and second (excluding MindGYM) in Table 1, outperforming other baselines.
+Notably, Open-O1(400) not only obtains the highest quality score among baselines, but also shows consistent Overall score improvements across all models compared to the raw version.
+MindGYM further surpass these baselines in quality: Qwen2.5-VL-7B reaches 0.93, while Qwen2.5-VL-32B achieves 0.98.
+These results suggest that larger LLMs not only produce better answers, but also generate higher-quality synthetic data.
+This trend holds across models and highlights the value of scaling for data generation.
+Lower Variance.
+Beyond high average quality, we observe a key phenomenon: as the parameter scale of the LLM used for data synthesis increases, the variance of quality scores decreases.
+For instance, Qwen2.5-VL-32B and InternVL3-38B exhibit low standard deviations (0.031 and 0.039, respectively), in contrast to baselines like MMEvol-DvQA, which shows a much higher variance (0.11) and ranks among the worst-performing in Table 1.
+Even within the baseline group, Open-O1(400) and MMEvol-SciQA, both with relatively lower variances, also achieve stronger Overall scores, typically outperforming their raw counterparts.
+Interestingly, although MMEvol-SciQA has slightly lower variance than Open-O1(400), its downstream performance lags behind, for example, showing negative Overall gain under Qwen2.5-VL-32B.
+This can be attributed to its lower mean quality score, reinforcing that both high average quality and low variance are critical to data effectiveness.
+The low-variance patterns observed in larger models suggest their superior ability not only to generate high-quality content but also to suppress outliers and noise—an essential property for stable and robust fine-tuning.
+These findings underscore a key insight: consistency matters as much as mean quality in building high-impact synthetic datasets.
+Richer Semantics.
+We next analyze linguistic properties that reflect structural complexity and task grounding: the average number of action verbs and the count of non-independent entities derived from dependency trees.
+MindGYM samples show significantly higher values in both metrics.
+For instance, Qwen2.5-VL-32B samples contain 14.14 action verbs and 2.14 non-independent entities on average, while LIMO and MMEvol-SciQA yield only 7.71/2.02 and 2.31/1.85, respectively.
+Meanwhile, MindGYM maintains a compact format, averaging 133–215 tokens and 92–110 characters per sample across models.
+This balance of rich semantics and concise form enhances model understanding while improving supervision efficiency and signal clarity.
+These results suggest that MindGYM prompts are both procedurally rich and semantically dense, which facilitates stronger self-consistency in instruction following.
+GPT-based Thinking Quality Scoring.
+We establish a two-dimensional evaluation framework leveraging GPT-4 as an expert scorer.
+The protocol assesses: (1) thinking Depth via derivation complexity, and (2) thinking Breadth by solution diversity.
+Each metric operates under a standardized scoring system, with detailed specifications provided in Appendix B.2.5.
+Comparative evaluations against four baselines are conducted on the MathVision dataset as a benchmark.
+The experimental results documented in Table 3 demonstrate MindGYM’s consistent superiority across both evaluation axes, particularly achieving 15.8% average improvements in thinking depth over the raw model and other competitive baselines.
 4.4 Ablation studies
-To validate the efficacy of our approach, we conduct systematic module-wise utility analysis through targeted ablation studies.
-GPT-based Thinking Quality Scoring.
-Datasets Depth Breadth Avg
-MindGYM/ Raw 10.2% 19.4% 14.8% $\uparrow$
-MindGYM/ LIMO 1.63% 23.8% 12.7% $\uparrow$
-MindGYM/ Open-O1 (400) 1.88% 23.7% 12.8% $\uparrow$
-MindGYM/ MMEvol-DVQA 8.41% 37.0% 22.7% $\uparrow$
-Overall 5.53% $\uparrow$ 26.0% $\uparrow$ 15.8% $\uparrow$
-Table 2: The results of win rates in terms of relative improvements over baselines, using GPT-4 as a scorer.
-We establish a two-dimensional evaluation framework leveraging GPT-4 as an expert scorer.
-The protocol assesses: (1) Reasoning Depth via derivation complexity, and (2) Reasoning Breadth by solution diversity.
-Each metric operates under a standardized scoring system, with detailed specifications provided in Appendix B.5.
-Comparative evaluations against four baselines are conducted on the MathVision dataset as a benchmark.
-The experimental outcomes documented in Table 2 demonstrate MindGYM’ consistent superiority across both evaluation axes, particularly achieving 15.8% average improvements in reasoning depth over the raw model and other competitive baselines.
+To validate the efficacy of our approach, we conduct systematic module-wise utility analysis through targeted ablation studies on Qwen2.5-VL-7Bqwen25vl .
+Ablation of Stream of Consciousness.
+Table 4: Ablation study results of removing stream of consciousness (w/o SC), utilizing English in data synthesis (Syn-EN), changing the order of our fine-tuning (w/o OF) steps and Relation Balanced (with RB), complete results detailed in Appendix D.
 Datasets MM-Avg Text-Avg Avg
@@
 MindGYM-Text Syn-EN 52.77 60.30 56.53 $\downarrow$
-MindGYM-Text w/o CF 52.20 60.83 56.51 $\downarrow$
+MindGYM-Text w/o OF 52.20 60.83 56.51 $\downarrow$
 MindGYM-Text with RB 52.20 60.83 56.51 $\downarrow$
-MindGYM-OKVQA 53.70 61.09 57.40
-MindGYM-OKVQA w/o SC 52.85 61.30 57.08 $\downarrow$
-MindGYM-OKVQA Syn-EN 53.14 60.59 56.87 $\downarrow$
-MindGYM-OKVQA with RB 53.33 60.60 56.97 $\downarrow$
-MindGYM-SciQA 53.11 61.46 57.29
-MindGYM-SciQA w/o SC 53.20 60.63 56.91 $\downarrow$
-MindGYM-SciQA Syn-EN 52.25 60.88 56.56 $\downarrow$
-MindGYM-SciQA with RB 53.46 60.27 56.87 $\downarrow$
-Table 3: Ablation study results of removing stream of consciousness (w/o SC), utilizing English in data synthesis (Syn-EN), changing the order of Curriculum Fine-tuning (w/o CF) steps and Relation Balanced (with RB), complete results detailed in Appendix C.
-Ablation of Stream of Consciousness.
-We conduct the ablation study by removing the stream of consciousness module from our prompt while maintaining the same curriculum fine-tuning method for training and evaluation.
-The experimental results, as shown in Table 3, indicate that the removal of the stream of consciousness module leads to a decrease across all Avg scores, with reductions ranging from 0.32 to 1.48 points.
-This decline highlights the significance of the stream of consciousness module in our approach, underscoring its essential role in enhancing the depth of reasoning and overall performance of the model.
-Curriculum Fine-tuning Order.
-To verify the effectiveness of our Curriculum Fine-tuning module, we mix and shuffle the first three datasets ($\mathcal{D}^{1}_{\text{train}}$, $\mathcal{D}^{2}_{\text{train}}$, $\mathcal{D}^{3}_{\text{train}}$) and train the model on text data, subsequently using $\mathcal{D}^{4}_{\text{train}}$ to constrain the answers’ generation pattern.
-This experimental setup allows us to systematically examine whether disrupting the curriculum order impacts model performance.
-The experimental results shown in Table 3 under the w/o CF condition reveal a performance degradation of 1.33 points, highlighting the critical impact of the Curriculum Fine-tuning order on model performance.
+We conduct the ablation study by removing the stream of consciousness module from our prompt while maintaining the same fine-tuning paradigm for training and evaluation.
+The experimental results, as shown in Table 4, indicate that the removal of the stream of consciousness module leads to a decrease across all Avg scores, with reduction of 1.48 points.
+This decline highlights the significance of the stream of consciousness module in our approach, underscoring its essential role in enhancing the depth of thinking and overall performance of the model.
+The Usage of Dataset.
+To verify the effectiveness of our fine-tuning module, we mix and shuffle the first three phase and train the LLM on text data, subsequently using the data of autonomous solving phase to constrain the answers’ generation pattern.
+This experimental setup allows us to systematically examine whether disrupting the fine-tuning order impacts model performance.
+The experimental results shown in Table 4 under the w/o OF condition reveal a performance degradation of 1.33 points, highlighting the critical impact of the our utilized order on the synthesis data.
 Linguistic Difference.
-Based on the selected Qwen series [4] model, we utilize Chinese as the language for data synthesis.
+Based on the selected Qwen qwen25vl , we utilize Chinese as the language for data synthesis.
 Additionally, we explore the effects of using English for synthesis, as shown in Table 4 under the Syn-EN condition.
-The results indicate varying degrees of performance degradation when using English, with a decrease of up to 1.31 points in the MindGYM-Text setting.
-Beyond the inherent characteristics of Qwen, we hypothesize that the higher information density of Chinese makes it more beneficial for the model’s reasoning learning.
+The results indicate the performance degradation when using English, with a decrease of 1.31 points.
+Beyond the inherent characteristics of Qwen, we hypothesize that the higher information density of Chinese makes it beneficial for the model’s thinking learning.
 Relationship Distribution Balanced.
 To verify the effectiveness of allowing the model to autonomously select relationship categories, we balance the number of entries generated for each category, totaling 400 entries.
-The experimental results shown in Table 3 under the with RB condition indicate that enforcing category balance leads to a decline up to 1.33 in overall model performance.
-This finding suggests that the autonomous selection of relationships by the model reflects its cognitive understanding.
+The experimental results shown in Table 4 under the with RB condition indicate that enforcing category balance leads to a decline of 1.33 in overall performance.
+This finding suggests that the autonomous selection of relationships by LLM reflects its cognitive understanding.
 5 Conclusion & Future Works
-This work introduces MindGYM, a framework that enhances vision-language reasoning via synthetic self-challenging questions.
-By self-challenging multi-step reasoning data and employing structured curriculum training, MindGYM achieves strong performance gains with small data size and reduced computational costs, outperforming prior methods in reasoning depth and breadth.
-Future directions include extending cognitive relationships to dynamic visual scenarios [46]; exploring the application of MindGYM in specific domains like medicine [27] and finance; developing adaptive scoring operators [6] for question complexity; and integrating agents [13] for data correctness verification in generalized RL environments.
-Appendix A Implementation of Experiments
-A.1 Utilized Seed Visual Datasets
-For the multimodal domain, we leverage widely-adopted scientific and reasoning datasets as visual source material for question synthesis, specifically including ScienceQA and OK-VQA.
-  • ScienceQA [22], the first large-scale multimodal dataset that annotates lectures and explanations for the answers.
-  • OK-VQA [25], a dataset for visual question answering that requires methods which can draw upon outside knowledge to answer questions.
-A.2 Benchmarks
+We present MindGYM, a cognitively guided framework for synthesizing self-challenging vision-language data.
+By injecting high-level reasoning signals and composing multi-hop tasks, MindGYM enables models to acquire structured thinking with minimal data and computation.
+Our synthetic data achieves higher quality and lower variance, leading to strong gains across reasoning benchmarks.
+Future directions include extending cognitive relationships to dynamic visual scenarios zhou2024humanvbenchexploringhumancentricvideo ; exploring the application of MindGYM in specific domains like medicine MURTAZA2023100546 and finance; developing adaptive scoring operators dj2 for question complexity; and integrating agents as for data correctness verification in generalized RL environments.
+Appendix
+Our appendix is organized as follows:
+  • Sec. A: The details of experimental setup, including the benchmarks (Sec. A.1) and the training & evaluation details (Sec. A.2).
+  • Sec. B: The implementation details of MindGYM, including the description of meta-topics (Sec. B.1) and all prompts utilized in this paper (Sec. B.2).
+  • Sec.
+C: Our exploration on multimodal data synthesis with MindGYM, including our proposed method on multimodal data (Sec.
+C.1), the experimental setup (Sec.
+C.2), experimental results (Sec.
+C.3) and the comprehensive analysis of the synthesis data (Sec.
+C.4).
+  • Sec. D: The ablation stuides, including the full experimental results.
+Appendix A Detailed Experimental Setup
+A.1 Benchmarks
 To evaluate the models’ reasoning capabilities and performance in both multimodal and text comprehension, we adopt the methodology used by major advanced open-source VLM Qwen2.5-VL-7B [3], selecting the following widely utilized reasoning evaluation datasets.
@@
 All experiments were conducted using this consistent setup to ensure the fairness of the experiments.
-A.3 Training and Evaluation Details
+A.2 Training and Evaluation Details
 Platform
@@
     • LIMO [39]: Challenges conventional assumptions in mathematical reasoning by demonstrating that models achieve superior performance with smaller quantities of high-quality training data.
-We reformat its ”question+solution+answer” structure into standard instruction pairs for training consistency.
+We reformat its "question+solution+answer" structure into standard instruction pairs for training consistency.
     • Open-O1 [32]: An open-source initiative to replicate the reasoning capabilities of proprietary models through curated SFT data for CoT activation.
@@
 For multimodal tasks, we implemented VLMEvalKit as the evaluation backend, while employing the Native framework for text-only evaluation scenarios.
-Appendix B Prompts
-B.1 Stream of Consciousness
+Appendix B Implementation Details of MindGYM
+B.1 Detailed Description of Meta-topics
+Unlike prior works focused on limited domain combinations [2], our approach emphasizes broad cognitive coverage:
+  • Mathematical Reasoning: Solving quantitative problems using numbers and formulas.
+  • Scientific Knowledge: Understanding natural laws and scientific principles.
+  • Logical Deduction: Forming conclusions through logical progression from premises.
+  • Technical Procedures: Following step-by-step instructions for practical tasks.
+  • Historical Events: Analyzing past events and their consequences.
+  • Ethical Considerations: Evaluating decisions through moral frameworks.
+  • Economic Principles: Studying resource allocation and valuation.
+  • Psychological Insights: Exploring human behavior and cognition.
+For purely textual contexts, we define three canonical combination types:
+  • Comparison: Identifying contrasts or similarities across entities or events;
+  • Bridging: Linking information across distant knowledge pieces via intermediate inference;
+  • Temporal: Reasoning across sequential or causal timelines.
+These types serve as explicit cues in $P_{\text{task}}^{\text{multi-hop}}$, guiding $\pi$ to adopt different reasoning schemata.
+B.2 Utilized Prompts
+B.2.1 Stream of Consciousness
 The protocol introduce in Section 3.2 is based on Thinking Claude, which is an innovative prompt engineering framework designed to enhance the reasoning capabilities of Claude AI models through simulated human-like cognitive processes.
@@
 We incorporate it into MindGYM’s data synthesis process to enhance the inference performance of the model [^3] https://github.com/richards199999/Thinking-Claude.
-B.2 Prompts for Background Generation
+B.2.2 Prompts for Background Generation
 On the text side, due to the low quality of sub-question generation without foundational documents, we create background documents for sub-question generation through iterative polling based on eight cognitive reasoning themes.
@@
 The document should be rich enough to support multi-hop reasoning in Chinese.
-B.3 Prompts for Seed Single-Hop Question Synthesis
+B.2.3 Prompts for Seed Single-Hop Question Synthesis
 These two prompts are used to synthesize sub-problems, Prompt 2 is used to synthesize text sub-problems, and Prompt 3 is used to synthesize images sub-problems.
@@
 Clearly state the relationship category that links all sub-questions in Chinese:
-B.4 Prompts for Challenging Multi-Hop QA Synthesis
+B.2.4 Prompts for Challenging Multi-Hop QA Synthesis
 These two prompts are used to synthesize multi-hop problems, Prompt 4 is used to synthesize multi-hop problems of text, and Prompt 5 is used to synthesize multi-hop problems of images.
@@
 Synthesize a multi-hop question and its answer based on the above sub-questions, original image, and original question in Chinese. Please start Qwen thinking and return the thinking process:
-B.5 Prompts for GPT-based Scorer
+B.2.5 Prompts for GPT-based Scorer
 This is a prompt scored with GPT4 to evaluate the scores of the two models in terms of depth and breadth.
@@
 3. Comparative analysis: use bullet points to list the main improvements/degradations
-Please return a JSON that strictly follows this format: ”accuracy_raw”: 0-2, ”accuracy_finetuned”: 0-2, ”depth_raw”: 0-3, ”depth_finetuned”: 0-3, ”comparison”: [”Point 1”, ”Point 2”]
-Appendix C Complete Results including Ablation Study
-We present the full experimental results in Table 4, comprehensively validating the effectiveness of our method compared to other baselines, as well as the robustness of our approach itself (ablation studies introduced in Section 4.3).
+Please return a JSON that strictly follows this format: "accuracy_raw": 0-2, "accuracy_finetuned": 0-2, "depth_raw": 0-3, "depth_finetuned": 0-3, "comparison": ["Point 1", "Point 2"]
+Appendix C Multimodal Data Synthesis Exploration
+C.1 Methodologies
+Anchoring Image Sources.
+For the $P_{context}$ component in text-image QA synthesis, we use existing images rather than generating images from scratch (as in the text-only case).
+This is because the existing SOTA VLMs are mostly only image-to-text generation (i.e., image understanding) and fail to do text-to-image generation at the same time.
+This decision arises because current SOTA VLMs predominantly specialize in image-to-text generation (i.e., image understanding) and lack dual capabilities for text-to-image generation constrained by their architectures.
+Without anchored visual references, the model risks hallucinatory reasoning about image elements (e.g., referencing “growth curves” might lead to incorrect inferences about unrelated scenarios, such as spatial-based reasoning tasks).
+Specifically, we randomly sample image-text pairs from scientific knowledge and general-domain QA datasets – ScienceQA [20] and OKVQA [22] – to serve as background context ($P_{\text{context}}$) for $\pi$.
+We emphasize alignment between visual and textual concepts within the generation instruction $P_{\text{task}}$, prompting $\pi$ to self-generate $k=5$ seed questions per inference.
+This enhances harmonized interactions between image and text modalities during synthesis.
+Multimodal Cognitive Combination Types.
+For $P_{\text{context}}$ containing both text and images, we introduce four multimodal combination types:
+  • Visual-Textual Alignment: Ensures consistency between visual data (e.g., diagrams, charts) and textual descriptions.
+  • Spatial Reasoning: Addresses spatial relationships or geometric arrangements depicted in visual content.
+  • Causal Inference: Requires deducing cause-effect relationships from combined textual and visual inputs.
+  • Contextual Synthesis: Demands integration of information across modalities (e.g., text, images, graphs) to form unified conclusions.
+Figure 2: An end-to-end example of MindGYM.
+For multimodal data, we generate five seed questions first, and make the model self-challenging itself via synthesizing multi-hop questions and multi-hop answers while preserving its internal thinking process.
+Figure 2 illustrates an example in which the original $P_{\text{context}}$ (including image and textual inputs) is first used to synthesize seed questions (e.g., “What ingredients did Kenji use?” or “What vitamins are in broccoli?”), as detailed in Section 3.2).
+These are then combined via the multi-hop synthesis protocol (outlined in Section 3.3) into more challenging questions that require cross-contextual reasoning (e.g., “How do Kenji’s lunch ingredients compare with Tracy’s in terms of vitamin content?” paired with bridging and temporal relationships).
+Critically, whereas the original QA focused on limited details (e.g., Kenji’s tomatoes and Tracy’s broccoli), our synthesized questions expand reasoning depth by incorporating broader contextual elements: sandwich compositions, citrus fruits, root vegetables, etc.
+The resulting multi-hop QAs demonstrate richer interconnections, broader scope, and higher cognitive demands on the solver.
+C.2 Experimental Setup
+While the main body of our work focuses exclusively on synthetic textual data, we also conduct preliminary explorations into multimodal data synthesis to demonstrate the extensibility of our approach.
+These experiments employ the same set of vision-language models—Qwen2.5-VL-7B-Instruct, Qwen2.5-VL-32B-Instruct, InternVL-8B, and InternVL-38B—for generating synthetic image-question-answer triplets.
+For the multimodal domain, we leverage widely-adopted scientific and reasoning datasets as visual source material for question synthesis, specifically including ScienceQA and OK-VQA.
+These datasets serve as visual grounding primitives, enabling the models to synthesize multimodal questions that align with real-world visual contexts.
+  • ScienceQA [20], the first large-scale multimodal dataset that annotates lectures and explanations for the answers.
+  • OK-VQA [22], a dataset for visual question answering that requires methods which can draw upon outside knowledge to answer questions.
+During training with image-text data, we update only the Aligner-layers, keeping both the ViT and LLM layers frozen to preserve pretrained representations and ensure stable adaptation.
+These results are not the main focus of the current study but demonstrate the extensibility of our method to multimodal contexts.
+Table 5: Performance of different models on various evaluation benchmarks.
+The number following the dataset name indicates its size in terms of the number of samples used for training.
+Different from Table 1, this table is about multimodal data synthesized by MindGYM.
+Text Eval Multimodal Eval
+Model Series Dataset GSM8K Math GPQA Text-Avg MMStar MathVista MathVision MM-Avg Avg
+[…]
+Openo1-sft (400) 84.31 69.00 29.70 61.00 63.87 69.70 23.66 53.08 57.04
+Openo1-sft (4k) 77.94 59.00 28.19 55.04 58.93 61.40 17.43 45.92 50.48
+[…]
+MMEvol-SciQA (106) 83.40 65.60 31.83 60.28 63.93 69.50 23.36 52.26 56.27
+MMEvol-DvQA (4k) 83.85 65.80 33.08 60.91 62.93 67.40 24.01 51.45 56.18
+[…]
+[…]
+[…]
+Openo1-sft (400) 95.38 81.20 50.51 75.70 68.80 73.40 36.18 59.46 67.58
+[…]
+[…]
+MMEvol-SciQA (106) 95.60 81.00 42.42 73.01 69.00 73.80 36.84 59.88 66.44
+MMEvol-DvQA (4k) 92.72 80.60 38.38 70.57 69.20 72.60 39.14 60.31 65.44
+[…]
+[…]
+[…]
+[…]
+[…]
+LIMO (817) 89.16 76.40 42.42 69.33 69.00 73.40 32.57 58.32 63.82
+[…]
+MMEvol-DvQA (4k) 88.55 74.60 40.40 67.85 68.47 71.60 32.24 57.44 62.65
+[…]
+[…]
+[…]
+[…]
+[…]
+LIMO (817) 89.16 72.60 47.47 69.74 72.40 72.80 36.51 60.57 65.16
+MMEvol-SciQA (106) 89.46 75.00 46.46 70.31 72.40 73.20 38.49 61.36 65.83
+MMEvol-DvQA (4k)} 89.92 74.00 46.46 70.13 70.93 71.70 30.92 57.85 63.99
+MindGYM-OKVQA(ours) 88.93 75.60 48.99 71.17 72.53 71.80 36.18 60.17 65.67
+[…]
+C.3 Experimental Results
+Cross-modal Transfer.
+MindGYM enables effective cross-modal reasoning capability sharing through bidirectional knowledge transfer.
+Remarkably, the text-specialized MindGYM-Text establishes new SOTA performance on MM-Avg, with the score of 54.42 that surpasses all baselines.
+Conversely, the multimodal-enhanced MindGYM-SciQA demonstrates unexpected supremacy in textual reasoning tasks, achieving an optimal Text-Avg score of 61.46.
+This bidirectional enhancement confirms the transferability of reasoning capabilities between textual and multimodal representations.
+Generalizability & Robustness.
+Consistent improvements emerge across diverse data sources.
+Our approach attains sub-optimal average performance when synthesizing data from both SciQA, showing an improvement of 0.45 over raw data, and OKVQA, with an improvement of 0.56 over raw data, demonstrating adaptability to various domain-specific reasoning tasks.
+This indicates that our method effectively enhances generalizability and robustness across diverse reasoning challenges.
+C.4 Data Analysis
+Table 6: Data-Juicer analysis results across different models. Different from Table 2, this table is about multimodal data synthesized by MindGYM.
+Model Dataset
+quality
+-mean
+quality
+-std
+action dependency token length
+baseline Openo1-sft (400) 0.96 0.091 8.05 2.00 77 274
+Openo1-sft (4k) 0.70 0.10 8.78 2.02 83 284
+LIMO (817) 0.87 0.10 7.71 2.02 101 322
+MMEvol-SciQA (106) 0.91 0.082 2.31 1.85 13.73 67
+MMEvol-DvQA (4k) 0.76 0.11 2.66 1.95 14.82 69
+Qwen2.5-VL-7B MindGYM-OKVQA 0.93 0.079 9.04 2.30 122 96
+MindGYM-SciQA 0.90 0.099 10.23 2.27 144 117
+Qwen2.5-VL-32B MindGYM-OKVQA 0.97 0.044 10.94 2.25 168 134
+MindGYM-SciQA 0.98 0.031 15.44 2.16 233 195
+InternVL-8B MindGYM-OKVQA 0.96 0.058 9.62 2.32 141 110
+MindGYM-SciQA 0.96 0.045 10.64 2.26 148 118
+InternVL-38B MindGYM-OKVQA 0.74 0.069 7.19 2.28 111 91
+MindGYM-SciQA 0.97 0.039 9.26 2.23 131 119
+We conduct a comparative analysis across baselines and our MindGYM-generated datasets to assess the impact of cognitively guided data on reasoning quality and complexity.
+As shown in Table 6, MindGYM datasets (OKVQA and SciQA) consistently yield higher reasoning quality scores, with Qwen2.5-VL-32B achieving the best performance (0.98 quality mean with 0.031 std on MindGYM-SciQA), significantly outperforming standard instruction-tuning sets such as Openo1-sft (0.70) and LIMO (0.87).
+In addition, MindGYM prompts models to generate substantially longer and more structured reasoning traces, as evidenced by the increase in action steps (up to 15.44) and dependency depth (avg.
+>2.2).
+This indicates that our cognitively annotated data effectively stimulates deeper reasoning behaviors without sacrificing generation stability.
+Notably, while token and input length grow with model capacity, the reasoning quality remains robust, suggesting that MindGYM scales well with model size and supports more abstract reasoning.
+In contrast, traditional instruction datasets lead to flatter, shorter outputs with limited cognitive structure.
+These results demonstrate the value of our data-centric approach in enhancing model reasoning ability and interpretability.
+Appendix D Comprehensive Ablation Studies
+We present the complete version of Table 4 in Table 7, which demonstrates the comprehensive ablation study results of our method.
+In addition to conducting ablation experiments on the text-based variant of our approach (as described in Appendix C), we also performed an ablation study on the multimodal data synthesis methodology.
+As shown in Table 7, the performance degradation observed in the OKVQA and SciQA ablation experiments compared to the full implementation of our method provides compelling evidence for two key aspects: 1) the effectiveness of our proposed framework, and 2) its cross-modal generalization capability across different data modalities.
+This empirical validation underscores the importance of our multimodal synthesis strategy in maintaining model performance across diverse vision-language tasks.
 Datasets Multimodal Eval Text-Only Eval Avg
@@
 MindGYM Syn-EN 63.93 69.70 24.67 52.77 84.00 67.20 29.70 60.30 56.53
-MindGYM w/o CF 64.00 69.90 22.69 52.20 84.15 67.00 31.33 60.83 56.51
+MindGYM w/o OF 64.00 69.90 22.69 52.20 84.15 67.00 31.33 60.83 56.51
 MindGYM w RB 64.00 70.10 22.70 52.20 84.15 67.00 31.33 60.83 56.51
@@
 MindGYM-SciQA w RB 64.20 70.20 25.99 53.46 83.85 65.00 31.96 60.27 56.87
-Table 4: Complete ablation study results of removing stream of consciousness (w/o SC), utilizing English in data synthesis (Syn-EN), changing the order of Curriculum Fine-tuning (w/o CF) steps and Relation Balanced (with RB).
-Appendix D Alternative VLMs Data Synthesis through MindGYM
-As demonstrated in Listing 1, our analysis using InternVL2.5-8B [5] reveals its generated outputs tend to exhibit chaotic patterns.
-We hypothesize this stems from dual factors: (i) constrained input token limits restricting contextual comprehension, and (ii) inherent deficiencies in cognitive reasoning capabilities.
-These limitations necessitated its exclusion from our comparative VLM experiments.
-Listing 1: Data Sample of MindGYM on InternVL2.5-8B.
-1{
-2 "background_document": "Background document for multi-hop reasoning.",
-3 "reasoning_category": "Mathematical Reasoning",
-4 "sub_questions": [
-5 "What is the optimal pricing strategy considering cost inflation?",
-6 "How will fixed and variable costs vary with seasonal demand?",
-7 "What are the impacts of geopolitical changes on supply chain?",
-8 "What role do regulatory changes play in pricing?",
-9 "How will budget allocation differ among different product lines?"
-10 ],
-11 "relationship_category": "Comparison",
-12 "multihop_question": "What steps need to be taken to optimize pricing strategies taking into account rising costs, shifting regulations, seasonal fluctuations, and differing product revenues?",
-13 "multihop_answer": "To maximize profit under increased costs, companies should reassess fixed and variable expenses, implement strategic cost savings wherever feasible, monitor regulatory developments closely, review existing budgets, allocate resources accordingly, update pricing schedules, communicate revisions internally consistently, measure effectivity quarterly.",
-[…]
-15}
+Table 7: Complete ablation study results of removing stream of consciousness (w/o SC), utilizing English in data synthesis (Syn-EN), changing the order of fine-tuning (w/o OF) steps and Relation Balanced (with RB).

### v3

@@
 Code and data are released to promote data-centric research into self-evolving foundation models driven by their internal reasoning capabilities.
-11footnotetext: Equal contribution.
+33footnotetext: FSIETP: Guangdong Provincial Key Laboratory of Fire Science and Intelligent Emergency Technology.
 1 Introduction
@@
 This highlights an important insight: not only does cognitively guided synthesis yield higher-quality data, but stability in quality—as measured by low variance—is critical for consistent fine-tuning.
+Compared with CoT wei2022chain and ToT yao2023tree baselines, our multi-step cognitive reasoning framework demonstrates superiority.
+We further observe that Chinese-synthesized data outperform English and mixed-language variants.
 Moreover, beyond the high-quality textual data synthesis presented in the main page, we also explore MindGYM’s scalability to multimodal settings in Appendix C, demonstrating its effectiveness and potential to both VLMs and large language models (LLMs).
@@
 Synthetic data generation has long been a tool for improving model reasoning.
-In language-only domains, question synthesis approaches have emphasized structure, difficulty scaling, and multi-step deduction, particularly in math or logic domains lewkowycz2022solving ; shao2024deepseekmath .
+In language-only domains, question synthesis approaches have emphasized structure, difficulty scaling, and multi-step deduction, particularly in math kuang2025atomic or logic domains lewkowycz2022solving ; shao2024deepseekmath .
 However, most existing methods either depend on rigid templates or domain-specific structures, which constrains their generalizability.
@@
 3 The Proposed MindGYM
-Figure 1: The proposed MindGYM framework incorporates a cognitively guided data synthesis pipeline with four stages: ① Context Generation, ② Single-Hop Question Synthesis, ③ Multi-Hop Composition with Thinking Trace, and ④ Structured Extraction.
+Figure 1: The proposed MindGYM framework incorporates a cognitively guided data synthesis pipeline with four stages: Context Generation, Single-Hop Question Synthesis, Multi-Hop Composition with Thinking Trace, and Structured Extraction.
 Starting from a meta-topic and guided by a shared cognitive objective, the model iteratively builds background context, atomic reasoning steps, composite questions with interpretable thinking traces, and final structured QA samples for downstream use.
@@
 These meta-topics span four core reasoning dimensions: quantitative, causal, temporal, and social-ethical, offering a comprehensive blueprint for cognitively rich question generation.
-Detailed meta-topics are provided in Appendix B.2.
+Detailed meta-topics are provided in Appendix A.2.
 Stage 1: Background Context Generation.
@@
 Reject Sampling for Diversity.
-To enhance diversity in seed question generation, MindGYM further applies semantic vectorization to textual components of data in the synthesis pool.
+To enhance diversity in seed question generation, MindGYM further applies semantic vectorization to textual components of data in the synthesis pool ling2025diversity .
 We use an empirically determined cosine similarity threshold to evaluate the overlap between newly generated and existing samples.
@@
 The task instruction $P_{\text{task}}^{\text{multi-hop}}$ specifies an abstract reasoning requirement (e.g., comparing entities, bridging concepts, or analyzing temporal relations), which guides the model to self-compose from prior seeds.
-Details of composition types are provided in Appendix B.
+Details of composition types are provided in Appendix A.
 Stage 4: Structured Extraction.
@@
 To comprehensively evaluate the performance of our proposed synthetic data method, we conduct extensive experimental validation on multiple representative evaluation sets.
-Full descriptions of datasets, benchmarks, implementation details, and baselines are provided in Appendix A.
+Full descriptions of datasets, benchmarks, implementation details, and baselines are provided in Appendix B.
 4.1 Experimental Settings
@@
 Data Efficiency.
-Despite using only 400 synthetic training samples, MindGYM achieves performance that not only matches but exceeds baselines trained with 10$\times$ more data.
-This includes outperforming Open-O1 (4k) in both textual and multimodal tasks.
-On MathVision, MindGYM reaches 28.62 while Open-O1 (4k) collapses to 17.43.
-This stark contrast underscores the high quality and utility of the data synthesized by MindGYM.
+[…]
 Cross-modal Enhancement.
@@
 The protocol assesses: (1) thinking Depth via derivation complexity, and (2) thinking Breadth by solution diversity.
-Each metric operates under a standardized scoring system, with detailed specifications provided in Appendix B.2.5.
+Each metric operates under a standardized scoring system, with detailed specifications provided in Appendix A.2.5.
 Comparative evaluations against four baselines are conducted on the MathVision dataset as a benchmark.
@@
 To validate the efficacy of our approach, we conduct systematic module-wise utility analysis through targeted ablation studies on Qwen2.5-VL-7Bqwen25vl .
-Ablation of Stream of Consciousness.
-Table 4: Ablation study results of removing stream of consciousness (w/o SC), utilizing English in data synthesis (Syn-EN), changing the order of our fine-tuning (w/o OF) steps and Relation Balanced (with RB), complete results detailed in Appendix D.
-Datasets MM-Avg Text-Avg Avg
+Impact of Structured Cognitive Thinking.
+Table 4: Ablation study results of removing structured cognitive thinking (w/o SC), utilizing English in data synthesis (Syn-EN), changing the order of our fine-tuning (w/o OF) steps and Relation Balanced (with RB), structured cognitive thinking with CoT and ToT synthes (CoT and ToT), complete results detailed in Appendix D.
+Datasets $M$-Avg Text-Avg Avg
 Raw 52.66 61.02 56.84
@@
 MindGYM-Text with RB 52.20 60.83 56.51 $\downarrow$
-We conduct the ablation study by removing the stream of consciousness module from our prompt while maintaining the same fine-tuning paradigm for training and evaluation.
-The experimental results, as shown in Table 4, indicate that the removal of the stream of consciousness module leads to a decrease across all Avg scores, with reduction of 1.48 points.
-This decline highlights the significance of the stream of consciousness module in our approach, underscoring its essential role in enhancing the depth of thinking and overall performance of the model.
+CoT 59.74 51.98 55.86 $\downarrow$
+ToT 59.78 52.07 55.92 $\downarrow$
+We investigate the role of the structured cognitive thinking (SC) module, which models multi-step reasoning behaviors such as planning, divergent exploration, and selective consolidation.
+This module is implemented through our multi-step synthesis scheme, where question and answer generation are explicitly decomposed into reasoning stages.
+To contrast with this design, we evaluate two widely used reasoning paradigms—Chain-of-Thought (CoT) and Tree-of-Thought (ToT)—which directly generate multi-hop questions without intermediate reasoning supervision.
+As shown in Table 4, MindGYM-Text achieves the highest average score of 57.84, outperforming both CoT (55.86) and ToT (55.92).
+This demonstrates that multi-step structured synthesis provides richer reasoning signals than direct single-pass generation.
+Moreover, removing SC (MindGYM-Text w/o SC) results in a performance drop to 56.36, confirming that structured cognitive guidance itself contributes to the improvement.
+Overall, these results suggest that multi-step, cognition-guided synthesis is more effective than direct CoT or ToT prompting for constructing high-quality reasoning data.
+Table 5: GPT-4o-based qualitative evaluation of synthesized data.
+logically flawed syntactically ambiguous correctness hallucination
+false 328 394 357 372
+true 71 5 42 27
+Effect of Data Quality and Filtering.
+We further analyze the quality and impact of our data through a controlled qualitative and filtering study.
+The dataset is evaluated by GPT-4o along four dimensions: logical consistency, linguistic clarity, factual correctness, and hallucination tendency.
+As shown in Table 5, over 89% of samples are judged factually correct, 93% non-hallucinatory, and fewer than 1% syntactically ambiguous, demonstrating the overall linguistic and logical soundness of the MindGYM synthesis process.
+Table 6: Impact of filtering on model performance.
+Model text-avg $M$-avg Overall Avg
+MindGYM (full) 61.27 54.42 57.84
+MindGYM (clean) 61.08 53.89 57.49
+To assess whether further cleaning benefits model training, we remove samples flagged as logically flawed or hallucinatory by GPT-4o and retrain the model on this “cleaned” subset.
+As presented in Table 6, the cleaned data lead to slightly lower performance compared to the full dataset (57.49 vs.
+57.84).
+This finding indicates that the effectiveness of MindGYM does not solely stem from factual accuracy, but rather from the rich reasoning trajectories embedded in the synthesized data.
+Even imperfect samples often encode meaningful intermediate reasoning steps—such as decomposition, hypothesis formation, and evidence evaluation—that contribute to more generalizable reasoning behavior.
+These results suggest that, for thinking-centric instruction tuning, maintaining reasoning diversity and cognitive structure is more beneficial than enforcing strict factual precision.
+Effect of Multimodal Synthesis and Vision Backbone Freezing.
+Table 7: Ablation on multimodal synthesis and vision backbone freezing.
+“CN” = Chinese text-only synthesis; “Flux” = additional text-to-image generation using Flux; “Frozen/Unfrozen” indicates whether the vision encoder is trainable.
+Setting text-avg $M$-avg Avg
+CN (text only, Frozen) 61.27 54.42 57.84
+Flux (image gen., Frozen) 60.91 52.68 56.80
+CN (text only, Unfrozen) 60.40 52.30 56.35
+We assess the influence of multimodal synthesis and vision backbone freezing.
+As shown in Table 7, the default Chinese text-only setting with a frozen encoder achieves the best overall average of 57.84.
+Introducing Flux-based image generation slightly reduces performance to 56.80, while unfreezing the vision backbone leads to a further drop to 56.35.
+These results indicate that additional multimodal synthesis and vision unfreezing introduce minor instability without clear benefit, validating our design choice of a text-dominant and frozen-encoder configuration for stable reasoning performance.
+For more information about Flux dataset synthesis, see the Appendix A.2.6.
+Linguistic Difference.
+Table 8: Cross-lingual comparison of MindGYM data synthesis.
+Setting text-avg $M$-avg Avg
+CN (Chinese only) 61.27 54.42 57.84
+MIX (CN + EN) 61.16 53.51 57.34
+EN (English only) 60.30 52.77 56.50
+Based on the selected Qwen qwen25vl , we examine the impact of synthesis language using three variants of MindGYM data: Chinese (CN), mixed Chinese–English (MIX; 50% Chinese and 50% English), and English (EN), as shown in Table 8.
+The Chinese-only configuration achieves the highest overall performance with an average score of 57.84 and is adopted as our default setting.
+In contrast, using English for synthesis leads to a noticeable performance drop of 1.34 points on average, while the mixed-language variant (MIX) also performs slightly worse than CN.
+These findings indicate that Chinese synthesis yields more effective reasoning supervision, likely benefiting from its higher information density and compact semantic structure, which align better with the thinking-centric nature of our data construction.
 The Usage of Dataset.
@@
 The experimental results shown in Table 4 under the w/o OF condition reveal a performance degradation of 1.33 points, highlighting the critical impact of the our utilized order on the synthesis data.
-Linguistic Difference.
-Based on the selected Qwen qwen25vl , we utilize Chinese as the language for data synthesis.
-Additionally, we explore the effects of using English for synthesis, as shown in Table 4 under the Syn-EN condition.
-The results indicate the performance degradation when using English, with a decrease of 1.31 points.
-Beyond the inherent characteristics of Qwen, we hypothesize that the higher information density of Chinese makes it beneficial for the model’s thinking learning.
 Relationship Distribution Balanced.
@@
 Future directions include extending cognitive relationships to dynamic visual scenarios zhou2024humanvbenchexploringhumancentricvideo ; exploring the application of MindGYM in specific domains like medicine MURTAZA2023100546 and finance; developing adaptive scoring operators dj2 for question complexity; and integrating agents as for data correctness verification in generalized RL environments.
+Acknowledgement
+This research is supported by Key-Area Research and Development Program of Guangdong Province, Granted No. 2024B1111060004.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: The abstract and introduction accurately summarize the paper’s main contributions and scope, which are consistently developed and supported throughout the paper (see Sec.
+3 and Sec.
+4.2).
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: Section 3.4 discusses the limitations of our approach and outlines potential directions for future work to extend our findings.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: This work focuses on developing a data synthesis method to generate high-quality training data, with primary emphasis on practical implementation rather than theoretical analysis.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: We propose the details of our data synthesis method in Sec. 3, with reproducibility details in Sec. 4.1 and Appendix B.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: All code and data are released to shed light on further research and applications.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We present the details of experimental setting in Sec. 4.1 and Appendix B.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34.
+Justification: Due to the considerable costs involved in training hundreds large language models, we run each experiment once and conduct extensive experiments across multiple benchmarks using different models to observe detailed results, which are reported in Sec.
+4.2 and  4.3.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: Experimental compute resources are provided in Appendix B.2.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: This research fully complies with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: Our societal impact and broader implications are discussed in Sec. 3.4
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: Our study involves neither high-risk data nor models.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: All models and datasets used in our experiments (Sec. 3 & Sec. 4.1) are publicly available, with proper attribution and adherence to their respective licenses.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: All code and data are open-sourced. See Section 3 for synthesis methods and Section 3.5 for usage instructions.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: Our data synthesis method employs self-supervised model generation, requiring no human intervention.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: No human subjects were involved in this study. All data were synthetically generated through self-supervised model training without human participation.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
 Appendix
-Our appendix is organized as follows:
-  • Sec. A: The details of experimental setup, including the benchmarks (Sec. A.1) and the training & evaluation details (Sec. A.2).
-  • Sec. B: The implementation details of MindGYM, including the description of meta-topics (Sec. B.1) and all prompts utilized in this paper (Sec. B.2).
-  • Sec.
-C: Our exploration on multimodal data synthesis with MindGYM, including our proposed method on multimodal data (Sec.
-C.1), the experimental setup (Sec.
-C.2), experimental results (Sec.
-C.3) and the comprehensive analysis of the synthesis data (Sec.
-C.4).
-  • Sec. D: The ablation stuides, including the full experimental results.
-Appendix A Detailed Experimental Setup
-A.1 Benchmarks
+Appendix A Implementation Details of MindGYM
+A.1 Detailed Description of Meta-topics
+Unlike prior works focused on limited domain combinations [2], our approach emphasizes broad cognitive coverage:
+  • Mathematical Reasoning: Solving quantitative problems using numbers and formulas.
+  • Scientific Knowledge: Understanding natural laws and scientific principles.
+  • Logical Deduction: Forming conclusions through logical progression from premises.
+  • Technical Procedures: Following step-by-step instructions for practical tasks.
+  • Historical Events: Analyzing past events and their consequences.
+  • Ethical Considerations: Evaluating decisions through moral frameworks.
+  • Economic Principles: Studying resource allocation and valuation.
+  • Psychological Insights: Exploring human behavior and cognition.
+For purely textual contexts, we define three canonical combination types:
+  • Comparison: Identifying contrasts or similarities across entities or events;
+  • Bridging: Linking information across distant knowledge pieces via intermediate inference;
+  • Temporal: Reasoning across sequential or causal timelines.
+These types serve as explicit cues in $P_{\text{task}}^{\text{multi-hop}}$, guiding $\pi$ to adopt different reasoning schemata.
+A.2 Utilized Prompts
+A.2.1 Stream of Consciousness
+The protocol introduce in Section 3.2 is based on Thinking Claude, which is an innovative prompt engineering framework designed to enhance the reasoning capabilities of Claude AI models through simulated human-like cognitive processes.
+This paradigm-shifting approach employs a multi-stage cognitive architecture featuring chain-of-thought prompting, self-reflective mechanisms, and structured problem decomposition.
+We incorporate it into MindGYM’s data synthesis process to enhance the inference performance of the model [^6] https://github.com/richards199999/Thinking-Claude.
+A.2.2 Prompts for Background Generation
+On the text side, due to the low quality of sub-question generation without foundational documents, we create background documents for sub-question generation through iterative polling based on eight cognitive reasoning themes.
+On the image side, since text-to-image generation models are still underdeveloped and produce low-quality results, we directly utilize existing reasoning and trial-based QA datasets.
+Write a background document of approximately 150-200 words.
+The document should describe a scenario or context that includes interconnected details, suitable for reasoning tasks.
+The document should focus on the reasoning category: {category}.
+The document should be rich enough to support multi-hop reasoning in Chinese.
+A.2.3 Prompts for Seed Single-Hop Question Synthesis
+These two prompts are used to synthesize sub-problems, Prompt 2 is used to synthesize text sub-problems, and Prompt 3 is used to synthesize images sub-problems.
+Based on the background document provided, generate up to 5 logically connected sub-questions.
+The relationship between these sub-questions should belong to a single category: - Bridging: requiring connecting facts or pieces of information from the document.
+- Comparison: involving comparing two or more elements described in the document.
+- Temporal: requiring reasoning about the order or timing of events.
+Clearly state the relationship category that links all sub-questions in Chinese:
+Based on the provided image, original question, and original answer, generate up to 5 logically connected sub-questions.
+The relationship between these sub-questions should belong to one of the following categories: - Visual-Textual Alignment: Requiring alignment between visual (e.g., images, charts) and textual information.
+- Spatial Reasoning: Involving spatial relationships or geometric layouts.
+- Causal Inference: Requiring reasoning about cause-and-effect relationships.
+- Contextual Synthesis: Requiring synthesis of information across multiple modalities (e.g., text, images, charts).
+Clearly state the relationship category that links all sub-questions in Chinese:
+A.2.4 Prompts for Challenging Multi-Hop QA Synthesis
+These two prompts are used to synthesize multi-hop problems, Prompt 4 is used to synthesize multi-hop problems of text, and Prompt 5 is used to synthesize multi-hop problems of images.
+Combine the sub-questions into a single, complex multi-hop question.
+The question should require reasoning across the sub-questions and synthesizing information from the background document.
+Then, provide a detailed answer to the multi-hop question, ensuring it is consistent with the background document and sub-questions.
+Synthesize a multi-hop question and its answer based on the above sub-questions and background document in Chinese.
+Please start Qwen thinking and return the thinking process:
+Combine the sub-questions into a single, complex multi-hop question.
+The question should require reasoning across the sub-questions and synthesizing information from the image and original question.
+Then, provide a detailed answer to the multi-hop question, ensuring it is consistent with the image and sub-questions.
+Synthesize a multi-hop question and its answer based on the above sub-questions, original image, and original question in Chinese.
+Please start Qwen thinking and return the thinking process:
+A.2.5 Prompts for GPT-based Scorer
+This is a prompt scored with GPT4 to evaluate the scores of the two models in terms of depth and breadth.
+As a mathematical problem solver, please strictly compare and analyze the predictions of the two models: [Problem description] {question}
+[Original model prediction] {pred_raw}
+[Fine-tuned model prediction] {pred_finetuned}
+Evaluation requirements: Give the scores of raw and finetuned respectively
+1. Reasoning depth score (0-3 points): - 3 points: multi-step derivation with verification - 2 points: complete derivation steps - 1 point: simple calculation steps - 0 points: no derivation process
+2.
+Reasoning breadth score (0-3): - 3 points: Explores multiple valid methods/angles, justifies optimal choice - 2 points: Mentions alternative approaches briefly but focuses on one - 1 points: Suggests another method without analysis - 0 points: Single approach with no alternatives considered
+3.
+Comparative analysis: use bullet points to list the main improvements/degradations Please return a JSON that strictly follows this format: "accuracy_raw": 0-2, "accuracy_finetuned": 0-2, "depth_raw": 0-3, "depth_finetuned": 0-3, "comparison": ["Point 1", "Point 2"]
+A.2.6 Prompts for Flux Data Sythesis
+We use the currently best text-to-image model Flux to generate images from text, and compare it with data based on ScienceQA and OKVQA as image anchors. The process involves two main steps:
+Step 1: Generating Structured Descriptions.
+We first produce concise textual descriptions that serve as prompts for image generation.
+These descriptions are categorized into eight distinct educational domains to ensure coverage of diverse reasoning scenarios:
+  • Geometry: basic geometric shapes with labeled lengths and angles.
+  • Physics: simple mechanics such as pulleys, inclined planes, or force diagrams.
+  • Chemistry: molecular structures or reaction schemes.
+  • Math Word Problem: real-world scenes involving objects and quantities.
+  • Logic Diagram: flowcharts or condition-based logical structures.
+  • Statistics Chart: visualizations such as bar charts, pie charts, or line graphs.
+  • Timeline or History Map: event timelines or migration maps.
+  • Circuit Diagram: basic electronic circuits with labeled components.
+Each category follows a unified generation prompt to ensure clarity and reliability: For each of the following categories, generate a simple and reliable prompt for a text-to-image generation model.
+The goal is to create educational images with accurate content, clear structure, and moderate visual complexity.
+The generated prompt should describe a plausible, textbook-style diagram with correct labels or values, suitable for multi-step reasoning (e.g.
+in math, science, or logic).
+The content should avoid errors, and not be too crowded or complex.
+Each prompt should be concise, standalone, and directly usable for image generation — do not include explanations or extra output.
+Use this category {category} and its description {description}:
+Step 2: Image Synthesis via Flux.
+The resulting category-specific prompts are then input into Flux, which generates corresponding images with consistent style and controlled complexity.
+Each synthesized image is paired with its associated textual reasoning data to form complete multimodal QA samples.
+This two-step pipeline—structured prompt generation followed by Flux-based synthesis—ensures both semantic precision and visual clarity, enabling the dataset to support reasoning across textual and visual modalities.
+Appendix B Detailed Experimental Setup
+B.1 Benchmarks
 To evaluate the models’ reasoning capabilities and performance in both multimodal and text comprehension, we adopt the methodology used by major advanced open-source VLM Qwen2.5-VL-7B [3], selecting the following widely utilized reasoning evaluation datasets.
@@
 Given the extensive evaluation tasks, employing the complete evaluation sets would incur substantial time expenditure.
-To expedite the evaluation process while ensuring fairness and accuracy, we adopt the $eval\_limit$ parameter from ms-swift, configuring GPQA to 300 samples, while retaining the original sample sizes for the remaining benchmark datasets.
+To expedite the evaluation process while ensuring fairness and accuracy, we adopt the eval_limiteval\_limit parameter from ms-swift, configuring GPQA to 300 samples, while retaining the original sample sizes for the remaining benchmark datasets.
 All experiments were conducted using this consistent setup to ensure the fairness of the experiments.
-A.2 Training and Evaluation Details
+B.2 Training and Evaluation Details
 Platform
@@
 For multimodal tasks, we implemented VLMEvalKit as the evaluation backend, while employing the Native framework for text-only evaluation scenarios.
-Appendix B Implementation Details of MindGYM
-B.1 Detailed Description of Meta-topics
-Unlike prior works focused on limited domain combinations [2], our approach emphasizes broad cognitive coverage:
-  • Mathematical Reasoning: Solving quantitative problems using numbers and formulas.
-  • Scientific Knowledge: Understanding natural laws and scientific principles.
-  • Logical Deduction: Forming conclusions through logical progression from premises.
-  • Technical Procedures: Following step-by-step instructions for practical tasks.
-  • Historical Events: Analyzing past events and their consequences.
-  • Ethical Considerations: Evaluating decisions through moral frameworks.
-  • Economic Principles: Studying resource allocation and valuation.
-  • Psychological Insights: Exploring human behavior and cognition.
-For purely textual contexts, we define three canonical combination types:
-  • Comparison: Identifying contrasts or similarities across entities or events;
-  • Bridging: Linking information across distant knowledge pieces via intermediate inference;
-  • Temporal: Reasoning across sequential or causal timelines.
-These types serve as explicit cues in $P_{\text{task}}^{\text{multi-hop}}$, guiding $\pi$ to adopt different reasoning schemata.
-B.2 Utilized Prompts
-B.2.1 Stream of Consciousness
-The protocol introduce in Section 3.2 is based on Thinking Claude, which is an innovative prompt engineering framework designed to enhance the reasoning capabilities of Claude AI models through simulated human-like cognitive processes.
-This paradigm-shifting approach employs a multi-stage cognitive architecture featuring chain-of-thought prompting, self-reflective mechanisms, and structured problem decomposition.
-We incorporate it into MindGYM’s data synthesis process to enhance the inference performance of the model [^3] https://github.com/richards199999/Thinking-Claude.
-B.2.2 Prompts for Background Generation
-On the text side, due to the low quality of sub-question generation without foundational documents, we create background documents for sub-question generation through iterative polling based on eight cognitive reasoning themes.
-On the image side, since text-to-image generation models are still underdeveloped and produce low-quality results, we directly utilize existing reasoning and trial-based QA datasets.
-[…]
-[…]
-The document should be rich enough to support multi-hop reasoning in Chinese.
-B.2.3 Prompts for Seed Single-Hop Question Synthesis
-These two prompts are used to synthesize sub-problems, Prompt 2 is used to synthesize text sub-problems, and Prompt 3 is used to synthesize images sub-problems.
-[…]
-[…]
-Based on the background document provided, generate up to 5 logically connected sub-questions. The relationship between these sub-questions should belong to a single category:
-- Bridging: requiring connecting facts or pieces of information from the document.
-- Comparison: involving comparing two or more elements described in the document.
-- Temporal: requiring reasoning about the order or timing of events.
-Clearly state the relationship category that links all sub-questions in Chinese:
-[…]
-[…]
-Based on the provided image, original question, and original answer, generate up to 5 logically connected sub-questions.
-The relationship between these sub-questions should belong to one of the following categories:
-- Visual-Textual Alignment: Requiring alignment between visual (e.g., images, charts) and textual information.
-- Spatial Reasoning: Involving spatial relationships or geometric layouts.
-- Causal Inference: Requiring reasoning about cause-and-effect relationships.
-- Contextual Synthesis: Requiring synthesis of information across multiple modalities (e.g., text, images, charts).
-Clearly state the relationship category that links all sub-questions in Chinese:
-B.2.4 Prompts for Challenging Multi-Hop QA Synthesis
-These two prompts are used to synthesize multi-hop problems, Prompt 4 is used to synthesize multi-hop problems of text, and Prompt 5 is used to synthesize multi-hop problems of images.
-[…]
-[…]
-Synthesize a multi-hop question and its answer based on the above sub-questions and background document in Chinese. Please start Qwen thinking and return the thinking process:
-[…]
-[…]
-The question should require reasoning across the sub-questions and synthesizing information from the image and original question.
-Then, provide a detailed answer to the multi-hop question, ensuring it is consistent with the image and sub-questions.
-Synthesize a multi-hop question and its answer based on the above sub-questions, original image, and original question in Chinese. Please start Qwen thinking and return the thinking process:
-B.2.5 Prompts for GPT-based Scorer
-This is a prompt scored with GPT4 to evaluate the scores of the two models in terms of depth and breadth.
-[…]
-[…]
-As a mathematical problem solver, please strictly compare and analyze the predictions of the two models:
-[Problem description]
-{question}
-[Original model prediction]
-{pred_raw}
-[Fine-tuned model prediction]
-{pred_finetuned}
-Evaluation requirements:
-Give the scores of raw and finetuned respectively
-1. Reasoning depth score (0-3 points):
-- 3 points: multi-step derivation with verification
-- 2 points: complete derivation steps
-- 1 point: simple calculation steps
-- 0 points: no derivation process
-2. Reasoning breadth score (0-3):
-- 3 points: Explores multiple valid methods/angles, justifies optimal choice
-- 2 points: Mentions alternative approaches briefly but focuses on one
-- 1 points: Suggests another method without analysis
-- 0 points: Single approach with no alternatives considered
-3. Comparative analysis: use bullet points to list the main improvements/degradations
-Please return a JSON that strictly follows this format: "accuracy_raw": 0-2, "accuracy_finetuned": 0-2, "depth_raw": 0-3, "depth_finetuned": 0-3, "comparison": ["Point 1", "Point 2"]
 Appendix C Multimodal Data Synthesis Exploration
@@
 Text Eval Multimodal Eval
-Model Series Dataset GSM8K Math GPQA Text-Avg MMStar MathVista MathVision MM-Avg Avg
+Model Series Dataset GSM8K Math GPQA Text-Avg MMStar MathVista MathVision $M$-Avg Avg
 […]
@@
 MMEvol-DvQA (4k) 92.72 80.60 38.38 70.57 69.20 72.60 39.14 60.31 65.44
-[…]
+MindGYM-OKVQA(ours) 95.53 82.20 $47 \ul$75.07 69.60 73.20 39.14 60.65 67.86
 […]
@@
 Appendix D Comprehensive Ablation Studies
-We present the complete version of Table 4 in Table 7, which demonstrates the comprehensive ablation study results of our method.
-In addition to conducting ablation experiments on the text-based variant of our approach (as described in Appendix C), we also performed an ablation study on the multimodal data synthesis methodology.
-As shown in Table 7, the performance degradation observed in the OKVQA and SciQA ablation experiments compared to the full implementation of our method provides compelling evidence for two key aspects: 1) the effectiveness of our proposed framework, and 2) its cross-modal generalization capability across different data modalities.
-This empirical validation underscores the importance of our multimodal synthesis strategy in maintaining model performance across diverse vision-language tasks.
 Datasets Multimodal Eval Text-Only Eval Avg
-MMStar MathVista MathVision MM-Avg GSM8K MATH GPQA Text-Avg
+MMStar MathVista MathVision $M$-Avg GSM8K MATH GPQA Text-Avg
 Baslines Raw 64.00 69.30 24.67 52.66 83.62 67.60 31.83 61.02 56.84
@@
 Table 11: Complete ablation study results of removing stream of consciousness (w/o SC), utilizing English in data synthesis (Syn-EN), changing the order of fine-tuning (w/o OF) steps and Relation Balanced (with RB).
+D.1 Complete Experimental Results
+We present the complete version of Table 4 in Table 11, which demonstrates the comprehensive ablation study results of our method.
+In addition to conducting ablation experiments on the text-based variant of our approach (as described in Appendix C), we also performed an ablation study on the multimodal data synthesis methodology.
+As shown in Table 11, the performance degradation observed in the OKVQA and SciQA ablation experiments compared to the full implementation of our method provides compelling evidence for two key aspects: 1) the effectiveness of our proposed framework, and 2) its cross-modal generalization capability across different data modalities.
+This empirical validation underscores the importance of our multimodal synthesis strategy in maintaining model performance across diverse vision-language tasks.
+D.2 Comparison with Recent Proprietary Systems
+Table 12: MathVision leaderboard comparison.
+Model Score
+GPT-4.5 47.3
+Gemini-2 Flash 41.3
+Qwen2.5-VL-72B-MindGym 39.47
+Kimi k1.5 38.6
+Qwen2.5-VL-72B (baseline) 38.10
+Claude 3.5 Sonnet 37.99
+Kimi-VL-A3B-Thinking 36.8
+To situate MindGym’s data-driven improvements in context, we compare the fine-tuned Qwen2.5-VL-72B-MindGym model with leading proprietary reasoning systems on the MathVision leaderboard.
+While MindGym is not a new architecture but a synthesis framework, it delivers a consistent +1.37% gain over the original Qwen2.5-VL model, narrowing the gap to top commercial models.
+Closed-source systems such as o3 and o4 remain inaccessible for fine-tuning, but MindGym effectively enhances open-source models under comparable evaluation settings.
+D.3 Comparison with MathFusion
+MindGYM is conceptually related to MathFusion [29] in that both approaches leverage synthetic data to improve reasoning capabilities.
+However, MindGYM extends this idea by integrating multimodal and cognitively structured synthesis, covering a wider range of domains beyond purely mathematical tasks.
+In particular, the Thinking Claude framework embeds structured reasoning trajectories into both questions and answers, providing richer supervision for reasoning-aware fine-tuning.
+To illustrate the difference, we compared a MathFusion-style synthesis with MindGYM on the same evaluation benchmarks:
+Table 13: Comparison between MathFusion-style synthesis and MindGYM.
+Method GSM8K Math GPQA text-avg MMStar MathVista MathVision $M$-avg Overall Avg
+MathFusion-style 84.31 69.1 29.0 60.80 61.43 67.2 24.97 51.20 56.00
+MindGYM 84.08 68.4 31.33 61.27 64.33 70.3 28.62 54.42 57.84
+As shown in Table 13, MindGYM outperforms the MathFusion-style baseline on multimodal and general reasoning tasks, demonstrating the effectiveness of combining cognitive structure with multimodal synthesis.